### v1

# Community detection for Contexual-LSBM: Theoretical limitation on misclassfication ratio and effecient algorithm

Abstract
The integration of both network information and node attribute information has recently gained significant attention in the context of community recovery problems. In this work, we address the task of determining the optimal classification rate for the Label-SBM(LSBM) model with node attribute information and. Specifically, we derive the optimal lower bound, which is characterized by the Chernoff-Hellinger divergence for a general LSBM network model with Gaussian node attributes. Additionally, we highlight the connection between the divergence $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ in our model and those introduced in Yun and Proutiere (2016) and Lu and Zhou (2016). We also presents a consistent algorithm based on spectral method for the proposed aggreated latent factor model.
1 Introduction
The community recovery problem Girvan and Newman (2002); Newman and Leicht (2007); Handcock et al. (2007); Fortunato (2010); Karrer and Newman (2011) is one of the fundamental challenges in the field of network science and data analysis. It involves identifying and reconstructing the underlying community structure within a network or graph. Accurately recovering these communities is crucial for understanding the organization and functionality of complex systems, ranging from social networks and biological systems to communication networks and financial markets.
Among the various proposed models, the Stochastic Block Model (SBM) Holland et al. (1983) stands out as one of the most significant and extensively studied in recent decades. Given known parameters $n$ and $K$, according to the definition of the SBM, each node with index $i\in[n]$ belonging to the set of nodes $\mathcal{V}$ is independently assigned to one of the $[K]$ clusters with probability $\alpha_{k}$. The probability of observing an edge $A_{ij}$ between any two nodes $i,j\in[N]$ is characterized by $\mathbb{P}_{\sigma(i),\sigma(j)}$ for a given label indicator function $\sigma(\cdot):[N]\rightarrow[K]$.
In addition to the basic SBM, numerous variants have been proposed and investigated, such as weighted/valued SBM Jog and Loh (2015); Aicher et al. (2015); Kurihara et al. (2006); Nowicki and Snijders (2001); Xu et al. (2020), degree-corrected SBM Karrer and Newman (2011); Gao et al. (2018); Yan et al. (2014); Qin and Rohe (2013), overlaping/mixed-membership SBM Airoldi et al. (2008); Latouche et al. (2014); Fu et al. (2009); Xing et al. (2010), hypergraph SBM Ghoshdastidar and Dukkipati (2014, 2015, 2017); Zhang and Tan (2023); Cole and Zhu (2020); Nandy and Bhattacharya (2024); Wang (2023); Gaudio and Joshi (2023), latent space model Hoff et al. (2002); Rohe et al. (2011); Smith et al. (2019); Koren (2005); Christensen and Golino (2021); Bing et al. (2020); Rohe and Zeng (2023); Bing et al. (2023) and so on. Among the aformentioned variants, Label-SBM(LSBM) Heimlicher et al. (2012), an extension of the SBM in terms of the non-binary observed edges, has recently garnered significant attention due to its capability in modeling diverse interaction types. More specfically speaking, the observed edge $l$ between any two nodes $u,v$ is sampled from a set ${0,1,2\cdots,L}$, where the probability is conditioned on the clusters $\sigma(u),\sigma(v)$ and and the label $l$, denoted as $\mathbf{P}(\sigma(u),\sigma(v),l)$. Moreover, real-world applications necessitate the analysis of contextual networks, as side information such as node attributes can be instrumental in uncovering the identity of node labels within groups. Substantial research has been dedicated to this area, as seen in Zhang et al. (2016); Abbe et al. (2022); Braun et al. (2022); Yan and Sarkar (2021). However, to the best of our knowledge, rigorous analysis of the integration of both types of information—Label-SBM and node attributes—has not yet been conducted.
In this paper, we investigate a model that combines the Label-Stochastic Block Model (LSBM) with contextual node attributes, referred to as the Contextual-Label-SBM (CLSBM). We assume a Gaussian mixture model for the contextual attribute vectors. Our primary focus is on analyzing the lower bound of the misclassification rate based on the model’s parameters. Specifically, let $s$ denote the expected number of misclassified nodes. Our objective is to establish the lower bound of the misclassification rate for the CLSBM when $s=o(n)$. While this result has been established for the LSBM model in (Yun and Proutiere, 2016, Theorem 1), our Theorem 1 is the first to extend this result to the CLSBM. Additionally, we aggregate topological and contextual information into a latent factor model and apply a spectral-based method to consistently estimate the community assignments for each node. The paper is organized as follows: In Section 2, we review related literature and outline our main contributions. In Section 3.1, we formalize the model definitions and discuss the assumptions necessary for our theoretical findings. Section 3.2 provides the theoretical lower bound for the number of misclassified nodes in the model and explores its connections with previous models. Finally, Section 3.3 introduces a spectral clustering algorithm for the proposed aggregated factor model and presents theoretical guarantees for its performance.
Notations
In this paper, bold lowercase letters, such as $\mathbf{x}$, represent vectors, while bold uppercase letters, like $\mathbf{X}$, denote matrices and $3$-d arrays. For a matrix $\mathbf{X}$, $\mathbf{X}_{ij}$ indicates the element located at the $i$-th row and $j$-th column, with $\mathbf{X}_{i\cdot}$ and $\mathbf{X}_{\cdot j}$ representing the $i$-th row and $j$-th column of $\mathbf{X}$, respectively. Similarly, for a $3$-d array $\mathbf{X}$, we use $\mathbf{X}_{i,j,k}$ to indicate the element at indices $i,j,k$ for each dimension. Occasionally, $\mathbf{x}_{j}$ is used as a simplified notation for $\mathbf{X}_{\cdot j}$.
The operators of probability and expectation are denoted by $\mathbb{P}$ and $\mathbb{E}$, respectively. For any integer $K\in\mathbb{N}$, the set $\{1,2,3,\cdots,K\}$ is represented by $[K]$. The notation $\left\|\mathbf{x}\right\|_{q}$, $\left\|\mathbf{X}\right\|_{F}$, $\left\|\mathbf{X}\right\|_{op}$ represent the $\ell_{p}$ norm of a vector $\mathbf{x}$, Frobenius norm and operoator norm of matrix $\mathbf{X}$ respectively. Here we use $\mathbf{1}(\cdot)$ to display the indicator function. For two sequences $a_{n}$ and $b_{n}$, the notation $a_{n}\lesssim b_{n}$ implies $a_{n}\leq Cb_{n}$ for some constant $C$.
2 Related work and main contribution
Jog and Loh (2015) was the first to investigate the threshold of exact recovery for the labeled/weighted SBM model by extending the results from Zhang and Zhou (2016). Additionally, Yun and Proutiere (2016) derived the optimal misclassification proportion for the LSBM and introduced a two-stage algorithm to achieve the information-theoretic limit. Besides, the contextual SBM, a generalization of SBM that incorporates additive node attributes to provide further information for identifying community clusteres, has been explored in various studies Braun et al. (2022); Deshpande et al. (2018); Dreveton et al. (2024); Abbe et al. (2022). However, to the best of our knowledge, the optimal classification rate for models with general non-binary labeled edges combined with independent arbitrary node attributes still remains unexplored. Our Theorem 1 shows that the lower bound for the expected misclasification number of nodes (as defined in Eq. 2) is $n\exp\left(-(1+\epsilon_{n})D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})n\right)$ for some sequence $\epsilon_{n}$ approaching zero. Here the key quantity $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$(defined in (4)) is the divergence defined based on the minimal value for the sum of two KL divergences, representing the topological and attribute information respectively. Furthermore, Remark 1 reveals a connection between the minimal value of sum of KL divergences over constrained sets and the Chernoff divergence, extending the results of (Yun and Proutiere, 2016, Claim $4$) to our settings and consequently prove that the aformentioned lower bound indeed match the optimal classfication rate in (Label)SBM(Zhang and Zhou (2016); Yun and Proutiere (2016)) and Gaussian mixture model(Lu and Zhou (2016)). One of our main technical contribution stems in the analysis for the solve of the problem (4). It is worth noting that, while the optimal solution for $D_{A}$ has been investigated in Yun and Proutiere (2016), the impact of the newly introduced term $D_{X}$ on the main results remains unclear. The additional term introduces further challenges, as it requires solving an optimization problem in function spaces, which is significantly more complex. In addition to establishing the theoretical lower bound, we also apply spectral methods to the proposed aggregated latent factor model, achieving a polynomial error rate. While this result is not optimal, the output of the spectral method can serve as a valuable initialization for further refinement in a second step, e.g see Gao et al. (2018); Yun and Proutiere (2016); Braun et al. (2022).
3 Model description and main result
3.1 Model definition
We primarily follow the notation introduced in Yun and Proutiere (2016). Each node is independently assigned to a cluster $\mathcal{V}_{k}$ with probability $\alpha_{k}$, where $k\in[K]$, and this assignment is independent of the node number $n$. Each edge $A_{i,j}\in\mathcal{V}^{2}$ is labeled with $l\in\{0,1,\cdots,L\}$, where the labeling probability depends on $i$, $j$, and $l$, and is denoted as $\mathbf{P}(\sigma(i),\sigma(j),l)$. Here, $\mathbf{P}\in\mathbb{R}^{K\times K\times(L+1)}$, and $\sigma(\cdot)$ is the node cluster indicator function. Here we summerize the above in the following definition of LSBM model as follows:
Definition 1 (Label Stochastic Block Model).
Let $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$ be an $3$-d tensor and we write $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}\sim LSBM\left(L,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$ if $\forall l\in\{0,1,\cdots,L\}$ $\mathbf{A}_{l}\in\left\{0,1\right\}^{n\times n}$ is symmertical and its upper triangular has independent elements such that
$\mathbb{P}(\mathbf{A}_{l}(i,j)=1)=\mathbf{P}(\sigma(i),\sigma(j),l),~{}\sum_{l=0}^{L}\mathbf{A}_{l}(i,j)=1,~{}\textit{and}\quad\mathbb{P}(\sigma(i)=k)=\boldsymbol{\alpha}_{k},$ (1)
$\forall i\neq j\in[n],\,k\in[K]$. Here we assume $\mathbf{A}_{l}(i,i)=0$ for all $i\in[K],l\in[L]$ to remove the self loop of the graph, $\boldsymbol{\alpha}$ is a constant vector independent of $n$ and $\sigma(\cdot):[n]\rightarrow[K]$ is the cluster assignment indicator function.
It is worth mentioning here that both $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$ and any third dimension of $\mathbf{P}$, such as $\mathbf{P}_{ij\cdot}\in\mathbb{R}_{+}^{(L+1)}$, must correspond to a probability distribution.
In addition to topological information, contextual attributes are also incorporated into the final model. We assume that each node $i$ is equipped with an attribute vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$, and the network $\mathbf{A}$ is considered independent of the sets of attribute vectors $\{\mathbf{x}_{i}\}$ given the label function $\sigma(\cdot)$. This setting was initially proposed in Deshpande et al. (2018) and has since been discussed in subsequent works Binkiewicz et al. (2017); Yan and Sarkar (2021); Weng and Feng (2016).
Definition 2 (Contexual Label Stochastic Block Model).
Let $d\geq 2$, $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$, $\boldsymbol{\mu}\in\mathbb{R}^{d\times K}$ ,and we write ($\left(\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}},\left\{\mathbf{x}_{i}\right\}_{i\in[n]}\right)\sim CLSBM(L,d,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot))$, if $\{\mathbf{A}_{l}\}_{l\in\{0,1,\cdots,L\}}\sim LSBM\left(n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$, and $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}\sim\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$.
Here we assume $\left\{\mathbf{A}_{l}\right\}_{l\in[L]}$ and $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$ are independently generated given the parameters respectively. Assume there exist a assignment vector fucntion $\bar{\sigma}(\cdot)$ to indicate the assignment for each node and this $\bar{\sigma}(\cdot)$ is estimated from the observed labels $\left\{\mathbf{A}_{l}\right\}_{l\in[L]\bigcup\{0\}]}$ and attribute $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$. In the next section, we provide the information theoretical lower-bound for the misclassification rate between estimated $\bar{\sigma}(\cdot)$ and ground truth $\sigma(\cdot)$ such that
$\bar{s}:=\inf_{\begin{subarray}{c}\pi\in S_{n}\\ \bar{\sigma}\in\mathcal{F}_{a}\end{subarray}}~{}\mathbb{E}\left[\sum_{i\in n}\mathbf{1}_{\{(\bar{\sigma}(\pi(i)))\neq\sigma(i))\}}\right].$ (2)
Here, $S_{n}$ represents the set of all permutations of $[K]$, and $\mathcal{F}_{a}$ denotes the set of all assignment functions $\sigma:[N]\to[K]$. Before presenting the main results, we first outline our assumptions.
Assumption 1.
$\exists~{}\eta_{1},\eta_{2}>0$ such that
$\forall l\in\{0,1,\cdots,L\},~{}\forall i,j,k\in[K],\quad\frac{\mathbf{P}(i,j,l)}{\mathbf{P}(i,k,l)}\leq\eta_{1},\quad\textit{and},\quad\left\|\boldsymbol{\mu}_{i}\right\|_{2}\leq\eta_{2}.$ (3)
The assumption mentioned earlier arises from the limitations of our proof techniques, and we believe it could potentially be eliminated with improved proof mechanisms. Notably, the assumption regarding $\mathbf{P}$ originates from Yun and Proutiere (2016), while the assumption for $\boldsymbol{\mu}$ is introduced to bound the second-order moments of the pseudolikelihood estimator for the perturbed model in the proof of Theorem 1. It is also worth mentioning that similar assumptions on $\boldsymbol{\mu}$ are commonly found in previous literature, such as Abbe et al. (2022).
3.2 Lower bound on number of misclassfication nodes
In this section we provide our main result. We start with the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$:
$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}):=\min_{k_{1}\neq k_{1}\in[K]}\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}_{+}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})+\frac{1}{n}D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X})$ (4)
where
$D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A}):=\max\left\{\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{1})\right),\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{2})\right)\right\},$
[…]
$D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X}):=\max\biggl{\{}\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{1})\right),\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{2})\right)\biggr{\}}.$ (5)
[…]
Theorem 1.
Denote $\bar{p}:=\max_{i,j,l\geq 1}\mathbf{P}(i,j,l)$, grant $\bar{p}=\omega(1/n)$, $\bar{p}=o(1)$ and 1. One has
lim $\liminf\frac{nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})}{\log(n/s)}\geq 1$ (6)
if $\eta_{2}=o(n)$. Here we assume $\limsup_{n\rightarrow\infty}\frac{\bar{s}}{s}\leq 1$. Here $\bar{s}$ is defined in (2).
Proof.
See Appendix A. ∎
Theorem 1 indeed implies that the number of misclassfied nodes for the proposed CLSBM model is above the order $n\exp(-nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}))$. The following remark indeed builds the connections between the Chernoff information $I$ proposed in (Dreveton et al., 2024, equation $3.4$) and $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ in (4)
Remark 1.
When $\bar{p}=o(1)$, for C-LSBM model, one has
$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$
$=\min_{k_{1}\neq k_{2}}\max_{t\in[0,1]}\sum_{k}^{K}\alpha_{k}\biggl{[}\sum_{l=1}^{L}\left((1-t)\mathbf{P}(k,k_{1},l)+t\mathbf{P}(k,k_{2},l)-\mathbf{P}^{1-t}(k,k_{1},l)\mathbf{P}^{t}(k,k_{2},l)\right)+\frac{t(1-t)}{2n}\left\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}\biggr{]}.$ (7)
Proof.
See Appendix B. ∎
From Remark 1, if the contextual attribute is missing such that $\boldsymbol{\mu}=0$, the divergence $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ reduces to the CH-divergence, as stated in Yun and Proutiere (2016). On the other hand, when the topological information is missing (e.g., $\mathbf{P}(k,k_{1},l)=\mathbf{P}(k,k_{2},l)$ for all $k,k_{1},k_{2},l$), we have $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})=\frac{1}{8n}\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\|_{2}^{2}$, which corresponds to the Gaussian mixture model, as shown in (Lu and Zhou, 2016, Theorem 3.3).
3.3 Spectral clustering for latent factor model
In this section, we extend the proposed CLSBM to a semiparametric model and demonstrate that the spectral method can consistently estimate the proposed estimator $s(\hat{\sigma})$ defined in (10), given the adjacency matrices $\{\mathbf{A}_{l}\}_{l\in\{0,1,2,\dots,L\}}$ and node attribute vectors $\{\mathbf{x}_{i}\}_{i\in[n]}$. Specifically, we first outline the construction of the estimator by combining the topological information $\{\mathbf{A}_{l}\}_{l\in\{0,1,2,\dots,L\}}$ with a Gaussian mixture model and provide an intuitive explanation of the spectral method at the population level in Section 3.3.1. Subsequently, in the following section, we formalize the spectral clustering method in Section 3.3.2 and establish the consistency of the proposed method, as proven in Lemma 1, to demonstrate its effectiveness.
3.3.1 Aggregated latent factor model
The latent factor model demonstrates its capability in revealing hidden low-dimensional structures within observed high-dimensional data and has been widely applied in various domains, including PCA, Gaussian mixtures, Nonnegative Matrix Factorization, and Sparse Factor Analysis. This provides an intuitive basis for combining both the topological model and Gaussian mixture into a shared latent space, as both the adjacency matrix and the observed Gaussian mixture share the same column space spanned by the ground truth assignment matrix $\mathbf{Z}(\sigma)\in\mathbb{R}^{n\times k}$, where $\mathbf{z}_{i\cdot}:=\mathbf{e}_{\sigma(i)}^{T}$. To elaborate, we begin by defining our aggregator factor matrix $\mathbf{S}$: given weights $\{w_{l}\}_{l\in[L]}$ independently and identically sampled from a uniform distribution over $[0,1]$, we define
$\mathbf{S}:=\sum_{l=1}^{L}w_{l}\mathbf{A}^{(l)}+\frac{1}{n}\mathbf{X}\mathbf{X}^{T}.$ (8)
where $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents the collection of $n$ observed Gaussian mixture data. The weights $\{w_{l}\}_{l\in[L]}$ are used to identify signals for different labels $l\in[L]$, and the factor $1/n$ balances the signal-to-noise ratio (SNR) between the graph signal and the Gaussian signal, as motivated by Remark 1. Let $\mathbf{P}_{s}:=\frac{1}{L}\sum_{l=1}^{L}\mathbf{P}_{\cdot,\cdot,l}\in\mathbb{R}^{K\times K}$. At the population level, we have:
$\mathbb{E}\left[\mathbf{S}\right]$ $=\mathcal{H}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})+\frac{1}{n}\mathbf{Z}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\mathbf{Z}^{T}$
$=\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}-\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})+\frac{1}{n}\mathbf{Z}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\mathbf{Z}^{T}$
$=\underbrace{\mathbf{Z}\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)\mathbf{Z}^{T}}_{\mathbf{M}}-\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}).$ (9)
Here, $\mathcal{H}(\cdot)$ denotes the hollowing operation, which removes the diagonal elements of the matrix, and $\mathbf{Z}$ is shorthand for $\mathbf{Z}(\sigma)$. It is evident that the first term $\mathbf{M}$ in $\mathbb{E}\left[\mathbf{S}\right]$ is a symmetric low-rank factor model, which has been extensively studied in areas such as low-rank matrix decomposition, Factor Analysis, signal processing, and spectral clustering. By the definition of $\mathbf{M}$, it is clear that the optimal solution of the $k$-means algorithm applied to the columns/rows of $\mathbf{M}$ should recover the ground truth assignment $\mathbf{Z}$. In practice, however, instead of directly applying $k$-means to the columns of $\mathbf{M}$ (which has dimension $n$), spectral methods are often employed due to their computational advantages. Specifically, let $\mathbf{U}_{K}$ denote the top $K$ eigenvectors of $\mathbf{M}$. We apply $k$-means to the columns of $\mathbf{U}_{K}^{T}\mathbf{M}$. This is equivalent to applying $k$-means to the columns of $\mathbf{M}$ or $\mathbf{U}_{K}^{T}\mathbf{M}$ in terms of clustering outcomes, as $\mathbf{U}_{K}$ is an orthonormal matrix that preserves distances during projection. However, since neither $\mathbf{M}$ nor $\mathbf{U}_{K}$ is observed, even at the population level, we substitute $\mathbf{S}$ for $\mathbf{M}$ and use its top eigenvectors, denoted as $\bar{\mathbf{U}_{K}}$, in place of $\mathbf{U}_{K}$. Given that the second term in (3.3.1) is small in operator norm, i.e., $\|\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})\|_{\text{op}}=o(1)$, the misclassification rate of the $k$-means algorithm based on $\mathbf{S}$ depends on $\|\mathbf{S}-\mathbb{E}(\mathbf{S})\|_{F}$, which can be bounded using concentration inequalities. We summarize the spectral clustering procedure in Algorithm 1.
1:Node attributes $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$, adjacent matrix $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}$, and number of clusters $K$
2:Cluster assignments for each data point
3:Step 1: Compute eigenvectors
4:Calculate the eigen decomposition of $\mathbf{S}$ as proposed in (8):
$\mathbf{S}=\sum_{i=1}^{n}\sigma_{i}\bar{\mathbf{u}_{i}}\bar{\mathbf{u}_{i}}^{T}.$
and denote $\bar{\mathbf{U}_{k}}\in\mathbb{R}^{n\times k}$ as the collection of the top $K$ eigenvectors.
5:Step 2: Apply $k$-means
6:Denote $\mathbf{Q}:=\bar{\mathbf{U}_{k}}^{T}\mathbf{S}\in\mathbb{R}^{k\times n}$ and denote $\bar{\sigma}(\cdot)$ as the solution to $K$-means such that:
$\bar{\sigma}(\cdot),\left\{\bar{\boldsymbol{\theta}}\right\}_{i\in[K]}=\operatorname*{argmin}_{\begin{subarray}{c}\sigma\in\mathcal{F}_{a},\\ \boldsymbol{\theta}_{i}\in\mathbb{R}^{k}\end{subarray}}\sum_{i\in[n]}\left\|\mathbf{Q}_{i}-\boldsymbol{\theta}_{\sigma(i)}\right\|_{2}^{2}.$
7:Return Cluster assignments function $\bar{\sigma}(\cdot)$.
Algorithm 1 Spectral Clustering Algorithm
3.3.2 Theoretical guarantee
Denote $s(\bar{\sigma})$ as the number of misclassified nodes for label function $\bar{\sigma}$ recovered from Algorithm 1 such that
$s(\bar{\sigma}):=\inf_{\pi\in S_{n}}~{}\mathbb{E}\left[\sum_{i\in n}\mathbf{1}_{\{(\bar{\sigma}(\pi(i)))\neq\sigma(i))\}}\right].$ (10)
The following lemma thus provides the upper bound of $s(\bar{\sigma})$.
Lemma 1.
Grant the assumptions in theorem 1 and further assume $\bar{p}=w(\log(n)/n)$, let $\bar{\sigma}$ be the solution of the spectral algorithm, we have
$s(\bar{\sigma})\lesssim\frac{K}{n\textit{SNR}}$ (11)
with high probability. Here $\textit{SNR}:=(1+o(1))\min_{k_{1}\neq k_{2}}\left\|\left(\mathbf{P}_{s}\right)_{k_{1}}-\left(\mathbf{P}_{s}\right)_{k_{2}}+\frac{1}{n}\left(\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right)^{T}\boldsymbol{\mu}\right\|_{2}^{2}$. Furthermore if $n\textit{SNR}=w(1)$ we obtain $s(\bar{\sigma})=o(1)$.
Proof.
See Appendix C
Lemma 10 demonstrates that the proposed indicator $s(\bar{\sigma})$ is upper bounded by the inverse of the defined (SNR), which corresponds to the minimum distance between any two columns in the matrix $\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}$. Rates of this nature have been extensively studied in the literature Lei and Rinaldo (2015); Löffler et al. (2021); Zhang and Zhou (2024) for spectral methods under SBM or Gaussian mixtures, respectively. However, our work is the first to provide theoretical guarantees for spectral methods under our proposed CLSBM. Despite achieving consistency, our analysis does not attain the optimal rate presented in Theorem 1, which is exponentially small. However, we beileive that this estimate could indeed play the role of a initializers for the potential refinement as seen in Yun and Proutiere (2016); Gao et al. (2018).
4 Conclusion
In this paper, we derive the theoretical lower bound for the misclassification rate in the Label-Stochastic Block Model with Gaussian node attributes. Our results also establish connections with previously introduced models, highlighting their broader applicability and usefulness. Beyond the theoretical lower bound, we propose a spectral method with theoretical guarantees for the proposed model, achieving a polynomial error rate in estimating community cluster assignments with high probability. In future work, we aim to explore the theoretical limitations of the spectral method to determine whether it can attain the optimal exponential rate.
Appendix A Proof of Theorem 1
Mimic the change of measure method in Yun and Proutiere (2016) we adopt it to the CLSBM model and its pertubted version in this section. Denote $\Phi$ and $\Psi$ as the parameter for the true parameter model(CLSBM) and one potential perturbed parameter model respectively. To be more specific, the true ground truth model $\Phi$ first generate the cluster for each node based on $\boldsymbol{\alpha}$ such that $\mathbb{P}(\sigma(i)=k)=\alpha_{k}$ for the $i$-th node indepedently. For the topological imformation, model $\Phi$ generate the observed label $l\in[L]$ for two different node $i$, $j$ based on the $\mathbf{P}(\sigma(i),\sigma(j),l)$. For the contexual information, $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$. We state the definition of $\Psi$ as follows:
Definition of the perturbed model $\Psi$
: Denote $(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star},q_{X}^{\star})$ as the solve for $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ such that
$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})=D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})+\frac{1}{n}D_{X}(k_{1}^{\star},k_{2}^{\star},q_{X}^{\star})$ (12)
The edge label generation mechanism for the topological information($\mathbf{A}$) in $\Psi$ is the same as in Yun and Proutiere (2016) by replace $\mathbf{q}$ with $\mathbf{q}_{\mathbf{A}}^{\star}$, here we state the generation of label for side information $\{\mathbf{x}_{i}\}_{i\in[n]}$ as follows: denote $i^{\star}=\operatorname*{argmin}_{\sigma(i)\in\{k_{1}^{\star},k_{2}^{\star}\}}i$.
  • When $i\neq i^{\star}$: $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$ which is same as in $\Phi$.
  • When $i=i^{\star}$: $\mathbf{x}_{i^{\star}}$ is sampled from the distribution whose pdf is $q_{\mathbf{X}}^{\star}$.
We denote the $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the pseudolikelihood ratio of the observed labels and contexual variables such that
$\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}:=\sum_{i\neq i^{\star}}^{n}\log\frac{\mathbf{q}_{A}^{\star}(\sigma(i),\mathbf{A}_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}.$ (13)
In the following two sections we thus provide the detailed proof of theorem 1, which follows the proof in Yun and Proutiere (2016). The proof consist of two parts, the first part constructs a stochastic model $\Psi$ and lower bound mis-classification node $\mathbb{E}_{\Phi}(\epsilon^{\pi}(n))$ by the expectation of the proposed log-likelihood ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ defined in (13); In the second part we further estimate the ratio upper bound of $\mathbb{E}_{\Psi}(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}})$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]\right)^{2}\right]$ respectively to finish the proof.
A.1 Part 1
Leverage Lemma 2 we assert there exist $(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star},q_{X}^{\star})$ such that
$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ $=\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{1}^{\star})\right)+\frac{1}{n}\operatorname{KL}_{X}\left(q_{X}^{\star},p_{X}(k_{1}^{\star})\right)$
$=\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{2}^{\star})\right)+\frac{1}{n}\operatorname{KL}_{X}\left(q_{X}^{\star},p_{X}(k_{2}^{\star})\right)$ (14)
Recall the definition of the log ratio Eq. 13:
$\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}:=\sum_{i\neq i^{\star}}^{n}\log\frac{\mathbf{q}_{A}^{\star}(\sigma(i),\mathbf{A}_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}.$ (15)
Mimic the approach described in Yun and Proutiere (2016), we could reach inequality $(17)$ in Yun and Proutiere (2016) adopted with our new defined log-ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the necessary condition for $\mathbb{E}[\epsilon(n)]\leq s$:
$\log(n/s)-\log(2/\alpha_{k_{1}})\leq\mathbb{E}_{\Psi}(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}})+\sqrt{\frac{4}{\alpha_{k_{1}}}\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]\right)^{2}\right]}.$ (16)
Then the next part finishes the proof by estiamting the bounds of $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}\right)^{2}\right]$ respectively.
A.2 Part 2
  • Upper-bound for $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$:
    Denote $L:=\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ and we obtain
    $\mathbb{E}_{\Psi}[L]$
    $:=\mathbb{E}_{\Psi}\left[\sum_{i\neq i^{\star}}^{n}\log\frac{q_{A}^{\star}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right].$
    $=\underbrace{\mathbb{E}_{\Psi}\sum_{i\neq i^{\star}}^{n}\log\frac{q_{A}^{\star}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}}_{\Gamma_{1}}+\underbrace{\mathbb{E}_{\Psi}\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}}_{\Gamma_{2}}$ (17)
    $\Gamma_{1}$ has been carefully studied in (Yun and Proutiere, 2016, Theorem 1) such that
    $\Gamma_{1}\leq(n+2\log^{2}(n)\log(\eta_{1}))D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star}).$
    Thus we will focus on $\Gamma_{2}$: By definition of $\Psi$ we have
    $\Gamma_{2}=D_{X}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{X}^{\star}).$
    Combine $\Gamma_{1}$ and $\Gamma_{2}$ together we have
    $\mathbb{E}_{\Psi}[L]\leq(n+2\log^{2}(n)\log(\eta_{1}))D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}).$
    Furthermore repeating the procedure in (A.2) leads to
    $\mathbb{E}_{\Psi}[L]\geq(n-2\log^{2}(n)\log(\eta_{1}))D(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu}).$
    Consequently we obtain
    $\left|\mathbb{E}_{\Psi}[L]-nD(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu})\right|\leq 2\log^{2}(n)\log(\eta)D(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu}).$ (18)
  • Upper bounding $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}\right)^{2}\right]$: Recall the notation of $L$ we obtain:
    $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}\right)^{2}\right]$
    $=\mathbb{E}_{\Psi}[\left(L-\mathbb{E}_{\Psi}[L]\right)^{2}]$
    $=\mathbb{E}_{\Psi}[L^{2}]-\left(\mathbb{E}_{\Psi}[L]\right)^{2}$
    $\leq\mathbb{E}_{\Psi}[L^{2}]-\left(n-2\log^{2}(n)\right)^{2}D^{2}(\boldsymbol{\alpha},\mathbf{p},\mu).$ (19)
    Then the following part remains to upper bound $\mathbb{E}_{\Psi}[L^{2}]$: Without loss of generaloity we assume $\sigma(i^{\star})=k_{1}$, by definition we have
    $\mathbb{E}_{\Psi}[L^{2}]$
    $=\mathbb{E}_{\Psi}[L^{2}|~{}\sigma(i^{\star})=k_{1}]$
    $=\underbrace{\mathbb{P}(i^{\star}\leq\log^{2}(n))\mathbb{E}_{\Psi}\biggl{[}L^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}\leq\log^{2}(n)\biggr{]}}_{\Gamma_{1}}+\underbrace{\mathbb{P}(i^{\star}>\log^{2}(n))\mathbb{E}_{\Psi}\biggl{[}L^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}>\log^{2}(n)\biggr{]}}_{\Gamma_{2}}$ (20)
    For $\Gamma_{2}$ we have
    $\Gamma_{2}$ $=\mathbb{P}(i^{\star}>\log^{2}(n))\mathbb{E}_{\Psi}\biggl{[}L^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}>\log^{2}(n)\biggr{]}$
    $\leq\frac{1}{n^{4}}\mathbb{E}_{\Psi}\left[\left(\sum_{i\neq i^{\star}}^{n}\log\frac{q_{A}^{\star}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}>\log^{2}(n)\right]$
    $=\frac{1}{n^{4}}\mathbb{E}_{\Psi}\biggl{[}\left(\sum_{i\neq i^{\star}}^{n}\log\frac{q_{A}^{\star}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}\right)^{2}+\left(\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}$
    $\quad\quad\quad+2\left(\sum_{i\neq i^{\star}}^{n}\log\frac{q_{A}^{\star}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}\right)\left(\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}>\log^{2}(n)\biggr{]}$
    $\leq\frac{1}{n^{4}}\left[n^{2}\log^{2}(\eta_{1})+\mathbb{E}_{x\sim q_{X}^{\star}}\left[\left\|\mathbf{x}-\boldsymbol{\mu}_{k_{1}}\right\|_{2}^{2}-\left\|\mathbf{x}-(1-t)\boldsymbol{\mu}_{k_{1}}-t\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}\right]^{2}+n\log(\eta_{1})\operatorname{KL}_{X}(q_{X}^{\star},p_{X}(\sigma(i^{\star})))\right]$
    $\lesssim\frac{1}{n^{4}}\left(n^{2}\log^{2}(\eta_{1})+t(1-t)\left(\eta_{2}^{4}+\eta_{2}^{2}\right)\right),$
    $\lesssim\frac{\log^{2}(\eta_{1})}{n^{2}}+\frac{\eta_{2}^{4}+\eta_{2}^{2}}{n^{4}}.$ (21)
    Here we use $\mathbb{P}(i^{\star}>\log^{2}(n))\leq n^{-4}$ in the second step, independence of $\{\mathbf{x}\}_{i\in[L]}$ and $\mathbf{A}$ in the third step, equation (33) and assumption 1 in the fourth steps and $10<t<$ 1 in the last step. To proceeds we then focus on $\Gamma_{1}$:
    $\Gamma_{1}$ $\leq\mathbb{E}_{\Psi}\biggl{[}L^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}\leq\log^{2}(n)\biggr{]}$
    $=\mathbb{E}_{\Psi}\biggl{[}\left(\sum_{i\neq i^{\star}}^{n}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}\leq\log^{2}(n)\biggr{]}$
    $=\mathbb{E}_{\Psi}\biggl{[}\left(\sum_{i>i^{\star}}^{n}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}\leq\log^{2}(n)\biggr{]}$
    $\quad+\mathbb{E}_{\Psi}\biggl{[}\left(\sum_{i<i^{\star}}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}\leq\log^{2}(n)\biggr{]}$
    $\quad+2\mathbb{E}_{\Psi}\biggl{[}\left(\sum_{i<i^{\star}}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}\right)\left(\sum_{i>i^{\star}}^{n}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)$
    $\quad\quad\quad\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}\leq\log^{2}(n)\biggr{]}$
    $\leq\underbrace{\mathbb{E}_{\Psi}\biggl{[}\left(\sum_{i>i^{\star}}^{n}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}\leq\log^{2}(n)\biggr{]}}_{\Gamma_{11}}$
    $\quad+\left(\log(n)\log(\eta_{1})+\eta_{2}\right)^{2}+2\left(\log(n)\log(\eta_{1})+\eta_{2}\right)D(\boldsymbol{\alpha},\mathbf{p},\mu).$ (22)
    Here we use the fact that any two nodes $i_{1}$ and $i_{2}$ such that $i_{1}\neq i_{2}\geq i^{\star}$ are muturally independent in the derivation. To upper bound $\Gamma_{11}$ we first have
    $\left(\sum_{i>i^{\star}}^{n}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}$
    $=\sum_{i,j>i^{\star}}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}\log\frac{q_{A}(\sigma(j),A_{j,i^{\star}})}{\mathbf{P}(\sigma(j),\sigma(i^{\star}),A_{j,i^{\star}})}+\left(\frac{n-i^{\star}}{n}\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}$
    $\quad+2\left(\frac{n-i^{\star}}{n}\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star})}\right)\sum_{i>i^{\star}}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}$ (23)
    We thus finish upper-bounding $\Gamma_{1}$ by analyzing $\Gamma_{11}$:
    $\Gamma_{11}$ $=\mathbb{E}_{\Psi}\biggl{[}\left(\sum_{i>i^{\star}}^{n}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}\leq\log^{2}(n)\biggr{]}$
    $=(n-i^{\star})^{2}D^{2}(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu})+\sum_{i>i^{\star}}\mathbb{E}\left[\log^{2}\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}\right]-\sum_{i>i^{\star}}\left(\mathbb{E}\left[\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}\right]\right)^{2}$
    $\leq(n-i^{\star})^{2}D^{2}(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu})+(n-i^{\star})\log(\eta_{1})D(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu}).$ (24)
    Plug (A.2) back to (A.2) and (A.2) sequentially we conclude
    $\mathbb{E}_{\Psi}[L^{2}]$ $\leq\frac{\log^{2}(\eta_{1})}{n^{2}}+\frac{\eta_{2}^{4}+\eta_{2}^{2}}{n^{4}}+\left(\log(n)\log(\eta_{1})+\eta_{2}\right)^{2}+2\left(\log(n)\log(\eta_{1})+\eta_{2}\right)D(\boldsymbol{\alpha},\mathbf{p},\mu)$
    $\quad+(n-i^{\star})^{2}D^{2}(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu})+(n-i^{\star})\log(\eta_{1})D(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu})$
    $\lesssim n^{2}D^{2}(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu})+\left(n\log(\eta_{1})+\eta_{2}\right)D(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu})$ (25)
    Further we end the proof by plugging above inequation back to (A.2):
    $\mathbb{E}_{\Psi}\left[\left(\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}-\mathbb{E}_{\Psi}[\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}]\right)^{2}\right]$
    $\leq\left(n+\log(n)\right)\left(\log(\eta_{1})D(\boldsymbol{\alpha},\mathbf{p},\mu)-D^{2}(\boldsymbol{\alpha},\mathbf{p},\mu)\right).$ (26)
Finally, we align (16), (A.2) and (A.2) together to conclude
lim $\liminf\frac{nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})}{\log(n/s)}\geq 1.$
Our proof then ends here.
Appendix B Proof of Remark 1
In this section we provide the proof of Remark 1:
Proof.
Recall the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$, pick any $k_{1}\neq k_{2}\in[K]$ and denote $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ such that
$D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}):=\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})+\frac{1}{n}D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X}).$ (27)
Here $D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})$ and $D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X})$ follow the definition in (3.2). Thus the proof remains to display
$D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$
$=\max_{t\in[0,1]}\sum_{k}^{K}\alpha_{k}\biggl{[}\sum_{l=1}^{L}\left((1-t)\mathbf{P}(k,k_{1},l)+t\mathbf{P}(k,k_{2},l)-\mathbf{P}^{1-t}(k,k_{1},l)\mathbf{P}^{t}(k,k_{2},l)\right)+\frac{t(1-t)}{2n}\left\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}\biggr{]}.$ (28)
We start the proof by studying the cosntrained version of (27): Notice that $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ actually reach the minimum of the following problem:
$\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{1})\right)+\frac{1}{n}\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{1})\right)$
[…]
$\quad>\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{2})\right)+\frac{1}{n}\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{2})\right).$ (29)
Thus we write the lagrange dual form of Appendix B such that
$\max_{t\geq 0}\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}L(\mathbf{q}_{A},q_{X},t)$ (30)
where
$L(\mathbf{q}_{A},q_{X},t)$
$=\underbrace{\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{1})\right)+t\left(\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{2})\right)-\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{1})\right)\right)}_{\Gamma_{1}(\mathbf{q}_{\mathbf{A}},t)}$
$\quad+\frac{1}{n}\left[\underbrace{\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{1})\right)+t\left(\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{2})\right)-\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{1})\right)\right)}_{\Gamma_{2}(q_{X},t)}\right]$ (31)
for some postive variable $t>0$. Notice that $\Gamma_{1}\left(\mathbf{q}_{A},t\right)$ has been completely studied in claim $4$ in Yun and Proutiere (2016), due to the independence of the $\mathbf{q}_{A},q_{X}$, the remaining part then focues on solving $\Gamma_{2}(q_{X},t)$.
Since $q_{X}$ represent a probability density function that lies in the function space $\mathcal{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Denote $\frac{\delta F(p)}{\delta p}(\mathbf{x})$ as the functional derivative of $F(p)$ to $p$, $\forall p\in\mathcal{F}$, by definition we have
$\frac{\delta\Gamma_{2}(q_{X},t)}{\delta q_{X}}(\mathbf{x})$
$=(1-t)\left[\log(q_{X}(\mathbf{x}))+1-\log(p_{X}(k_{1})(\mathbf{x}))\right]+t\left[\log(q_{X}(\mathbf{x}))+1-\log(p_{X}(k_{2})(\mathbf{x})\right].$ (32)
Here we use $\frac{\delta\operatorname{KL}(q,p)}{\delta q}(\mathbf{x})=\log(q(\mathbf{x}))+1-\log(p(\mathbf{x}))$ in the derivation. Denote $q_{X}^{\star}(x):=\operatorname*{argmin}_{q_{X}\in\mathcal{F}}\Gamma_{2}(q_{X},t)$ and let (B) equal to zero we have:
$\log(q_{X}^{\star}(\mathbf{x}))=(1-t)\log(p_{X}(k_{1})(\mathbf{x}))+t\log(p_{X}(k_{2})(\mathbf{x}))-1.$
due to the convexity of $\Gamma_{2}(q_{X},t)$. Notice that $p_{X}(k_{1})$ and $p_{X}(k_{2})$ represent the pdf for $\mathcal{N}(\boldsymbol{\mu}_{k_{1}},\mathbf{I}_{d})$ and $\mathcal{N}(\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ respectively, consequently we obtain
$q_{X}^{\star}(\mathbf{x})=C\exp(-\left\|\mathbf{x}-(1-t)\boldsymbol{\mu}_{k_{1}}-t\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}/2).$ (33)
after some calculation. Here $C$ is some constant that is irelevent with $\mathbf{x}$. We assert $C=(2\pi)^{-d/2}$ by utilzing $q_{X}$ is an probability density function.
Plug the expression of $q_{X}^{\star}$ back to $\Gamma_{2}(q_{X},t)$ defined in (B) gives:
$\Gamma_{2}(q_{X}^{\star},t)=\frac{t(1-t)}{2}\left\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}.$ (34)
after some algebra. Here we use Lemma 3 and the fact that $q_{X}^{\star}$ is the PDF for $\mathcal{N}((1-t)\boldsymbol{\mu}_{k_{1}}+t\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ in the derivation. We thus finish the proof by combining the expression of $\Gamma_{1}(\mathbf{q}_{A},t)$(see (Yun and Proutiere, 2016, Claim $4$)) and $\Gamma_{2}(q_{X}^{\star},t)$ together.
Appendix C Proof of Lemma 1
In this section we provide the proof of Lemma 1:
Proof.
Let $\mathbf{S}^{(k)}$ be the best rank-$k$ approximation of $\mathbf{S}$, define $\hat{\sigma}(\cdot)$ and $\{\mathbf{c}_{i}\}_{i\in[n]}$ such that
$\hat{\sigma}(\cdot),\{\hat{\mathbf{c}_{i}}\}_{i\in[K]}:=\operatorname*{argmin}_{\begin{subarray}{c}\sigma(\cdot)\in\mathcal{F},\\ \mathbf{c}_{i}\in\mathbb{R}^{n}\end{subarray}}\sum_{i\in[K]}\left\|\mathbf{S}^{(k)}_{i}-\mathbf{c}_{\sigma(i)}\right\|_{2}^{2}.$ (35)
By lemma 4 we have $\hat{\sigma}=\bar{\sigma}\circ\pi$ for some permutation $\pi$. Thus we could then focues on $\hat{\sigma}$ in the remaining part of the proof. We start by bounding $\left\|\mathbf{S}^{(k)}-\mathbf{M}\right\|_{op}$ for the $\mathbf{M}$ defined in (3.3.1):
$\left\|\mathbf{S}^{(k)}-\mathbf{M}\right\|_{op}$ $\leq\left\|\mathbf{S}^{(k)}-\mathbb{E}\left[\mathbf{S}\right]\right\|_{op}+\left\|\mathbb{E}\left[\mathbf{S}\right]-\mathbf{M}\right\|_{op}$
$\leq\left\|\mathbf{S}^{(k)}-\mathbf{S}\right\|_{op}+\left\|\mathbf{S}-\mathbb{E}\left[\mathbf{S}\right]\right\|_{op}+\left\|\mathbb{E}\left[\mathbf{S}\right]-\mathbf{M}\right\|_{op}$
$\leq 2\left\|\mathbf{S}-\mathbb{E}\left[\mathbf{S}\right]\right\|_{op}+\left\|\mathbb{E}\left[\mathbf{S}\right]-\mathbf{M}\right\|_{op}$
$\leq 2\left(\left\|\sum_{l=1}^{L}w^{(l)}\mathbf{A}^{(l)}-\mathbb{E}\left[\sum_{l=1}^{L}w^{(l)}\mathbf{A}^{(l)}\right]\right\|_{op}+\left\|\frac{1}{n}\mathbf{X}\mathbf{X}^{T}-\frac{1}{n}\mathbb{E}(\mathbf{X}\mathbf{X}^{T})\right\|_{op}\right)+1$
$\lesssim\sqrt{n}.$ (36)
with probability at least $1-cn^{-c^{\prime}}$. Here we use $\mathbf{S}^{(k)}$ is the best approximation of $\mathbf{S}$ in the third step, (Coja-Oghlan, 2010, Lemma 8.5) and (Löffler et al., 2021, Lemma B.1) in the last step. Notice that $\operatorname{rank}~{}(\mathbf{S}^{(k)}-\mathbf{M})\leq\operatorname{rank}(\mathbf{S}^{(k)})+\operatorname{rank}(\mathbf{M})\leq 2K$ we further have
$\left\|\mathbf{S}^{(k)}-\mathbf{M}\right\|_{F}\leq\sqrt{2k}\left\|\mathbf{S}^{(k)}-\mathbf{M}\right\|_{op}\lesssim\sqrt{Kn}.$ (37)
with high probability. On the other hand define $\hat{\mathbf{C}}$ such that $\hat{\mathbf{C}}_{i}:=\hat{\mathbf{c}}_{\hat{\sigma}(i)}$, consequently we have $\hat{\mathbf{C}}$ is the solve of (35) :
$\left\|\hat{\mathbf{C}}-\mathbf{S}^{(k)}\right\|_{F}\leq\left\|\mathbf{M}-\mathbf{S}^{(k)}\right\|_{F}.$ (38)
Consequently we have
$\left\|\hat{\mathbf{C}}-\mathbf{M}\right\|_{F}\leq\left\|\hat{\mathbf{C}}-\mathbf{S}^{(k)}\right\|_{F}+\left\|\mathbf{S}^{(k)}-\mathbf{M}\right\|_{F}\leq 2\left\|\mathbf{S}^{(k)}-\mathbf{M}\right\|_{F}\lesssim\sqrt{Kn}.$ (39)
with high probability. Here we use (37) in the last step. Define set $S_{m}(\delta)$ such that
$S_{m}(\delta):=\left\{i\in[n]~{}|~{}\left\|\hat{\mathbf{C}_{i}}-\mathbf{M}_{i}\right\|_{2}\geq\frac{\delta}{2}\right\}.$ (40)
where $\delta:=\min_{\mathbf{M}_{i}\neq\mathbf{M}_{j}}\left\|\mathbf{M}_{i}-\mathbf{M}_{j}\right\|_{2}$. It is easy to check any index $i\in[n]\setminus S_{m}(\delta)$ will not be misclassfied up to permutation. Thus we finish the proof by upper bound the size of $S_{m}(\delta)$:
$s(\hat{\sigma})\leq\textit{card}~{}(S_{m}(\delta))\leq\frac{4\left\|\hat{\mathbf{C}}-\mathbf{M}\right\|_{F}^{2}}{\delta^{2}}\lesssim\frac{kn}{\delta^{2}}.$ (41)
By the definition of $\delta$ we have
$\delta$ $=\min_{\sigma(i)\neq\sigma(j)}\left\|\mathbf{Z}\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)\mathbf{z}_{i}^{T}-\mathbf{Z}\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)\mathbf{z}_{j}^{T}\right\|_{2}$
$=\min_{\sigma(i)\neq\sigma(j)}\left\|\mathbf{Z}\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)_{\sigma(i)}-\mathbf{Z}\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)_{\sigma(j)}\right\|_{2}$
$=\min_{\sigma(i)\neq\sigma(j)}\left\|\mathbf{Z}\boldsymbol{\Sigma}^{-1}\boldsymbol{\Sigma}\left(\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)_{\sigma(i)}-\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)_{\sigma(j)}\right)\right\|_{2}$
$=\min_{\sigma(i)\neq\sigma(j)}\left\|\boldsymbol{\Sigma}\left(\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)_{\sigma(i)}-\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)_{\sigma(j)}\right)\right\|_{2}$ (42)
where $\boldsymbol{\Sigma}:=\operatorname{diag}([n_{1},n_{2},\cdots,n_{K}])=n\left(\operatorname{diag}(\boldsymbol{\alpha})+o(1)\mathbf{I}_{K}\right)$. Here we use the orthonormality for the columns of $\mathbf{Z}\boldsymbol{\Sigma}^{-1}$ in the last step. Thus we further obtain
$\delta=(1+o(1))n\min_{\sigma(i)\neq\sigma(j)}\left\|\left(\mathbf{P}_{s}\right)_{\sigma(i)}-\left(\mathbf{P}_{s}\right)_{\sigma(j)}+\frac{1}{n}\left(\boldsymbol{\mu}_{\sigma(i)}-\boldsymbol{\mu}_{\sigma(j)}\right)^{T}\boldsymbol{\mu}\right\|_{2}.$ (43)
Combine (41) and (43) together we finish the proof here. ∎
Appendix D Technical lemmas
In this section we provide some technical lemmas that is essential to our main results.
Lemma 2.
Recall the definition of $D_{A}$, $D_{X}$ and $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$, one has
$D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})=\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{1})\right)=\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{2})\right)$
[…]
Here $\operatorname{KL}(\cdot,~{}\cdot)$ represent the KL divergence and
$(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star},q_{X}^{\star}):=\operatorname*{argmin}\limits_{\begin{subarray}{c}k_{1}\neq k_{2}\\ \mathbf{q}_{A}\in\mathbb{R}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}D_{A}(k_{1},k_{2},\mathbf{q}_{A})+\frac{1}{n}D_{X}(k_{1},k_{2},q_{X})$ (45)
Proof.
Proof of first equation in (2):
Here we mimic the approach in Yun and Proutiere (2016) and finish the proof by contradiction: Recall the definition of $D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})$ we have
$D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})=\max\left\{\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{1})\right),\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{2})\right)\right\}.$
Without loss of generality, we assume
$D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})=\sum_{k=1}^{K}\boldsymbol{\alpha}_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{1}^{\star})\right)>\sum_{k=1}^{K}\boldsymbol{\alpha}_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{2}^{\star})\right).$
Thus there should exist a $k^{\prime}\in[K]$ such that
$\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k^{\prime}),\mathbf{P}(k^{\prime},k_{1}^{\star})\right)>\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k^{\prime}),\mathbf{P}(k^{\prime},k_{2}^{\star})\right).$
Leverage the positivity of the KL-divergence we assert $\mathbf{q}_{A}^{\star}(k^{\prime})\neq\mathbf{P}(k^{\prime},k_{1}^{\star})$ then there should exist a $\mathbf{q}_{A}^{\prime}(\cdot,~{}\cdot)\in\mathbb{R}^{K\times L}$ satisfy the following two conditions
$\begin{cases}\operatorname{KL}_{A}(\mathbf{q}_{A}^{\prime}(k),\mathbf{P}(k,k_{1}^{\star}))=\operatorname{KL}_{A}(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{1}^{\star}))&k\neq k^{\prime}\\ \operatorname{KL}_{A}(\mathbf{q}_{A}^{\prime}(k),\mathbf{P}(k,k_{1}^{\star}))\in[\operatorname{KL}_{A}(\mathbf{q}_{A}^{\star}(k^{\prime}),\mathbf{P}(k^{\prime},k_{1}^{\star}))-\varepsilon,\operatorname{KL}_{A}(\mathbf{q}_{A}^{\star}(k^{\prime}),\mathbf{P}(k^{\prime},k_{1}^{\star}))]&k=k^{\prime}\end{cases}$
and
$\begin{cases}\operatorname{KL}_{A}(\mathbf{q}_{A}^{\prime}(k),\mathbf{P}(k,k_{2}^{\star}))=\operatorname{KL}_{A}(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{2}^{\star}))&k\neq k^{\prime}\\ \operatorname{KL}_{A}(\mathbf{q}_{A}^{\prime}(k),\mathbf{P}(k,k_{2}^{\star}))\in[\operatorname{KL}_{A}(\mathbf{q}_{A}^{\star}(k^{\prime}),\mathbf{P}(k^{\prime},k_{2}^{\star})),\operatorname{KL}_{A}(\mathbf{q}_{A}^{\star}(k^{\prime}),\mathbf{P}(k^{\prime},k_{2}^{\star}))+\varepsilon]&k=k^{\prime}\end{cases}$
where $\varepsilon$ is an constant such that $\varepsilon\in[0,1/2\left(\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k^{\prime}),\mathbf{P}(k^{\prime},k_{1}^{\star})\right)-\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k^{\prime}),\mathbf{P}(k^{\prime},k_{2}^{\star})\right)\right)]$. Then we have
$D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})>\underbrace{\sum_{k=1}^{K}\boldsymbol{\alpha}_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\prime}(k),\mathbf{P}(k,k_{1}^{\star})\right)}_{:=D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\prime})}>\sum_{k=1}^{K}\boldsymbol{\alpha}_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\prime}(k),\mathbf{P}(k,k_{2}^{\star})\right).$
thus further leads to:
$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})=D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})+D_{X}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{X}^{\star})>D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\prime})+D_{X}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{X}^{\star}).$
which is contradicted by the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$.
Proof of second equation in (2):
The proof here is in spirits similar to the previous part via contradicting the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ by constructing a new distribution function $q_{X}^{\prime}$: Recall the definition of the $D_{X}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})$:
$D_{X}(k_{1}^{\star},k_{2}^{\star},q_{X}^{\star}):=\max\left\{\operatorname{KL}_{X}\left(q_{X}^{\star},p_{X}(k_{1}^{\star})\right),\operatorname{KL}_{X}\left(q_{X}^{\star},p_{X}(k_{2}^{\star})\right)\right\}.$
W.L.O.G we assume
$\operatorname{KL}_{X}\left(q_{X}^{\star}(k),p_{X}(k_{1}^{\star})\right)>\operatorname{KL}_{X}\left(q_{X}^{\star}(k),p_{X}(k_{2}^{\star})\right).$ (46)
Similarly to the first part, we could also build a new measure $q_{X}^{\prime}(\cdot,\cdot)$ such that
$D_{X}(k_{1}^{\star},k_{2}^{\star},q_{X}^{\star})>\underbrace{\operatorname{KL}_{X}\left(q_{X}^{\prime},p_{X}(k_{1}^{\star})\right)}_{:=D_{X}(k_{1}^{\star},k_{2}^{\star},q_{X}^{\prime})}>\operatorname{KL}_{X}\left(q_{X}^{\prime}(k),p_{X}(k_{2}^{\star})\right).$
thus further leads to:
$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})=D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})+D_{X}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{X}^{\star})>D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})+D_{X}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{X}^{\prime}).$
which is a contradiction against the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$. Until then our proof ends. ∎
Lemma 3 (KL divergence for two Gaussian distributions).
Consider two Gaussian probability distributions with unit variance, having means $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$, and probability density functions (pdfs) denoted as $p_{1}$ and $p_{2}$, respectively. It follows that one has
$\operatorname{KL}(p_{1},p_{2})=\frac{\left\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\right\|_{2}^{2}}{2}.$ (47)
Here $\operatorname{KL}(p,q):=\mathbb{E}_{p}\log(\frac{p}{q})$ represents the KL-divergence for two continuous probability distribution.
Proof.
By definition we have
$\operatorname{KL}(p_{1},p_{2})$ $=\mathbb{E}_{\mathbf{x}\sim p_{1}}\log(\frac{p_{1}(\mathbf{x})}{p_{2}(\mathbf{x})})$
$=\frac{1}{2}\mathbb{E}_{\mathbf{x}\sim p_{1}}\left(\left\|\mathbf{x}-\boldsymbol{\mu}_{2}\right\|_{2}^{2}-\left\|\mathbf{x}-\boldsymbol{\mu}_{1}\right\|_{2}^{2}\right)$
$=\frac{1}{2}\left(\left\|\boldsymbol{\mu}_{2}\right\|_{2}^{2}-\left\|\boldsymbol{\mu}_{1}\right\|_{2}^{2}\right)+\mathbb{E}_{x\sim p_{1}}\left(\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\right)^{T}\mathbf{x}$
$=\frac{1}{2}\left(\left\|\boldsymbol{\mu}_{2}\right\|_{2}^{2}-\left\|\boldsymbol{\mu}_{1}\right\|_{2}^{2}\right)+\left(\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\right)^{T}\boldsymbol{\mu}_{1}$
$=\frac{\left\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\right\|_{2}^{2}}{2}.$ (48)
Lemma 4.
Let $\mathbf{S}^{(k)}$ as the best rank-$k$ approximation of $\mathbf{S}$ proposed in algorithm 1, denote $\hat{\sigma}(\cdot)$ and $\{\mathbf{c}_{i}\}_{i\in[n]}$ such that:
$\hat{\sigma}(\cdot),\{\mathbf{c}_{i}\}_{i\in[n]}:=\operatorname*{argmin}_{\begin{subarray}{c}\hat{\sigma}(\cdot)\in\mathcal{F},\\ \mathbf{c}_{i}\in\mathbb{R}^{n}\end{subarray}}\sum_{i\in[n]}\left\|\mathbf{S}^{(k)}_{i}-\mathbf{c}_{\hat{\sigma}(i)}\right\|_{2}^{2}.$ (49)
Recall the definition of $\bar{\sigma}(\cdot)$ in (1) we have $\bar{\sigma}(\pi(i))=\hat{\sigma}(i)$ for some permutation $\pi(\cdot)$.
Proof.
By definition of $\bar{\sigma}(\cdot)$ and $\mathbf{Q}$ in Algorithm 1 we have
$\sum_{i\in[n]}\left\|\mathbf{Q}_{i}-\boldsymbol{\theta}_{\bar{\sigma}(i)}\right\|_{2}^{2}$ $=\sum_{i\in[n]}\left\|\mathbf{U}_{k}\mathbf{Q}^{(k)}_{i}-\mathbf{U}_{k}\boldsymbol{\theta}_{\bar{\sigma}(i)}\right\|_{2}^{2}$
$=\sum_{i\in[n]}\left\|\mathbf{S}^{(k)}_{i}-\mathbf{U}_{k}\boldsymbol{\theta}_{\bar{\sigma}(i)}\right\|_{2}^{2}$
thus there exist $\hat{\sigma(i)}$ such that $\mathbf{c}_{\hat{\sigma}(i)}:=\mathbf{U}_{k}\boldsymbol{\theta}_{\bar{\sigma}(i)}$. Here we use $\mathbf{U}_{k}$ is a orthogional matrix is the first step and definition of $\mathbf{S}^{(k)}$ in the second step. ∎

### v3

@@
-Community detection for Contexual-LSBM: Theoretical limitation on misclassfication ratio and effecient algorithm
+Community Detection for Contextual-LSBM: Theoretical Limitations of Misclassification Rate and Efficient Algorithms
 Abstract
-The integration of both network information and node attribute information has recently gained significant attention in the context of community recovery problems.
-In this work, we address the task of determining the optimal classification rate for the Label-SBM(LSBM) model with node attribute information and.
-Specifically, we derive the optimal lower bound, which is characterized by the Chernoff-Hellinger divergence for a general LSBM network model with Gaussian node attributes.
-Additionally, we highlight the connection between the divergence $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ in our model and those introduced in Yun and Proutiere (2016) and Lu and Zhou (2016).
-We also presents a consistent algorithm based on spectral method for the proposed aggreated latent factor model.
+The integration of network information and node attribute information has recently gained significant attention in the community detection literature.
+In this work, we consider community detection in the Contextual Labeled Stochastic Block Model (CLSBM), where the network follows an LSBM and node attributes follow a Gaussian Mixture Model (GMM).
+Our primary focus is the misclassification rate, which measures the expected number of nodes misclassified by community detection algorithms.
+We first establish a lower bound on the optimal misclassification rate that holds for any algorithm.
+When we specialize our setting to the LSBM (which preserves only network information) or the GMM (which preserves only node attribute information), our lower bound recovers prior results.
+Moreover, we present an efficient spectral-based algorithm tailored for the CLSBM and derive an upper bound on its misclassification rate.
+Although the algorithm does not attain the lower bound, it serves as a reliable starting point for designing more accurate community detection algorithms (as many algorithms use spectral method as an initial step, followed by refinement procedures to enhance accuracy).
 1 Introduction
-The community recovery problem Girvan and Newman (2002); Newman and Leicht (2007); Handcock et al.
-(2007); Fortunato (2010); Karrer and Newman (2011) is one of the fundamental challenges in the field of network science and data analysis.
-It involves identifying and reconstructing the underlying community structure within a network or graph.
-Accurately recovering these communities is crucial for understanding the organization and functionality of complex systems, ranging from social networks and biological systems to communication networks and financial markets.
+The community detection problem Girvan and Newman (2002); Newman and Leicht (2007); Fortunato (2010); Karrer and Newman (2011) is one of the fundamental challenges in the field of network science and data analysis.
+It focuses on identifying and reconstructing the underlying community structure within a network or graph.
+Accurately identifying these communities is crucial for understanding the structure and function of complex systems, ranging from social networks and biological systems to communication networks and financial markets.
 Among the various proposed models, the Stochastic Block Model (SBM) Holland et al.
 (1983) stands out as one of the most significant and extensively studied in recent decades.
-Given known parameters $n$ and $K$, according to the definition of the SBM, each node with index $i\in[n]$ belonging to the set of nodes $\mathcal{V}$ is independently assigned to one of the $[K]$ clusters with probability $\alpha_{k}$.
-The probability of observing an edge $A_{ij}$ between any two nodes $i,j\in[N]$ is characterized by $\mathbb{P}_{\sigma(i),\sigma(j)}$ for a given label indicator function $\sigma(\cdot):[N]\rightarrow[K]$.
-In addition to the basic SBM, numerous variants have been proposed and investigated, such as weighted/valued SBM Jog and Loh (2015); Aicher et al.
-(2015); Kurihara et al.
-(2006); Nowicki and Snijders (2001); Xu et al.
-(2020), degree-corrected SBM Karrer and Newman (2011); Gao et al.
+Given parameters $n$ and $K$, according to the definition of the SBM, each node with index $i\in[n]$ belonging to the set of nodes $\mathcal{V}$ is independently assigned to the $k$-th community (where $k\in[K]$) with probability $\alpha_{k}$.
+Denote $\sigma:[n]\rightarrow[K]$ as the ground truth indicator function such that $\sigma(i)=k$ if the node $i$ belongs to community $k$, thus the community detection problem is to infer the ground truth indicator function $\sigma(\cdot)$ from the observed graph (usually represented by an adjacency matrix $\mathbf{A}$).
+Here the probability of observing an edge $\mathbf{A}_{ij}$ between any two nodes $i,j\in[n]$ is characterized by $\mathbb{P}_{\sigma(i),\sigma(j)}$.
+In addition to the basic SBM, many variants have been proposed and investigated, such as weighted/labeled SBM Jog and Loh (2015); Aicher et al.
+(2015); Nowicki and Snijders (2001); Yun and Proutiere (2016), degree-corrected SBM Karrer and Newman (2011); Gao et al.
 (2018); Yan et al.
-(2014); Qin and Rohe (2013), overlaping/mixed-membership SBM Airoldi et al.
+(2014); Qin and Rohe (2013), overlapping/mixed-membership SBM Airoldi et al.
 (2008); Latouche et al.
@@
 (2009); Xing et al.
-(2010), hypergraph SBM Ghoshdastidar and Dukkipati (2014, 2015, 2017); Zhang and Tan (2023); Cole and Zhu (2020); Nandy and Bhattacharya (2024); Wang (2023); Gaudio and Joshi (2023), latent space model Hoff et al.
+(2010), hypergraph SBM Ghoshdastidar and Dukkipati (2014); Zhang and Tan (2023); Cole and Zhu (2020); Nandy and Bhattacharya (2024), latent space model Hoff et al.
 (2002); Rohe et al.
-(2011); Smith et al.
-(2019); Koren (2005); Christensen and Golino (2021); Bing et al.
+(2011); Bing et al.
 (2020); Rohe and Zeng (2023); Bing et al.
-(2023) and so on.
-Among the aformentioned variants, Label-SBM(LSBM) Heimlicher et al.
-(2012), an extension of the SBM in terms of the non-binary observed edges, has recently garnered significant attention due to its capability in modeling diverse interaction types.
-More specfically speaking, the observed edge $l$ between any two nodes $u,v$ is sampled from a set ${0,1,2\cdots,L}$, where the probability is conditioned on the clusters $\sigma(u),\sigma(v)$ and and the label $l$, denoted as $\mathbf{P}(\sigma(u),\sigma(v),l)$.
-Moreover, real-world applications necessitate the analysis of contextual networks, as side information such as node attributes can be instrumental in uncovering the identity of node labels within groups.
-Substantial research has been dedicated to this area, as seen in Zhang et al.
+(2023).
+Among the aforementioned variants, in this paper we focus on the labeled SBM (LSBM) Heimlicher et al.
+(2012); Yun and Proutiere (2016), an extension of the SBM in terms of the non-binary observed edges.
+Compared to the basic SBM, the non-binary edge formulation provides enhanced capability in modeling diverse types of interactions between any pair of nodes.
+In the community detection literature, many existing methods rely solely on topological information, such as adjacency or Laplacian matrices.
+However, real-world applications often require the analysis of contextual networks, as side information like node attributes can play a crucial role in uncovering node labels within communities.
+Significant research has been dedicated to this area, as evidenced by works such as Zhang et al.
 (2016); Abbe et al.
@@
 (2022); Yan and Sarkar (2021).
-However, to the best of our knowledge, rigorous analysis of the integration of both types of information—Label-SBM and node attributes—has not yet been conducted.
-In this paper, we investigate a model that combines the Label-Stochastic Block Model (LSBM) with contextual node attributes, referred to as the Contextual-Label-SBM (CLSBM).
-We assume a Gaussian mixture model for the contextual attribute vectors.
-Our primary focus is on analyzing the lower bound of the misclassification rate based on the model’s parameters.
-Specifically, let $s$ denote the expected number of misclassified nodes.
-Our objective is to establish the lower bound of the misclassification rate for the CLSBM when $s=o(n)$.
-While this result has been established for the LSBM model in (Yun and Proutiere, 2016, Theorem 1), our Theorem 1 is the first to extend this result to the CLSBM.
-Additionally, we aggregate topological and contextual information into a latent factor model and apply a spectral-based method to consistently estimate the community assignments for each node.
-The paper is organized as follows: In Section 2, we review related literature and outline our main contributions.
-In Section 3.1, we formalize the model definitions and discuss the assumptions necessary for our theoretical findings.
-Section 3.2 provides the theoretical lower bound for the number of misclassified nodes in the model and explores its connections with previous models.
-Finally, Section 3.3 introduces a spectral clustering algorithm for the proposed aggregated factor model and presents theoretical guarantees for its performance.
-Notations
-In this paper, bold lowercase letters, such as $\mathbf{x}$, represent vectors, while bold uppercase letters, like $\mathbf{X}$, denote matrices and $3$-d arrays.
+Specifically, Zhao et al.
+(2021); Sima et al.
+(2021); Abbe et al.
+(2022); Dreveton et al.
+(2024); Braun et al.
+(2022) investigate the threshold for exact recovery (equavilent to strong consistency Abbe (2018)) of the Contextual-SBM, with Zhao et al.
+(2021); Sima et al.
+(2021); Abbe et al.
+(2022) focusing on the special case of two communities and Dreveton et al.
+(2024); Braun et al.
+(2022) extending the results to the more general case of $K$ communities.
+While exact recovery is a valid performance metric that has received significant theoretical attention, it requires the number of misclassified nodes to be exactly zero, which is overly stringent and often impractical.
+In contrast, most real-world applications of community detection or node classification often prioritize weak consistency, where a small fraction of nodes is allowed to be misclassified.
+This aligns more closely with practical needs, where partial misclassification is acceptable and more realistic.
+Motivated by practical considerations, in this paper, we investigate a model that integrates the LSBM with contextual Gaussian node attributes, referred to as Contextual-LSBM (CLSBM).
+Our focus is on the relationship between the misclassification rate and the model parameters.
+Specifically, let $s$ represent the expected number of misclassified nodes.
+Our primary objective is to establish the lower bound of the misclassification rate for the CLSBM when $s=o(n)$.
+While this result has been established for the LSBM in (Yun and Proutiere, 2016, Theorem 1), our Theorem 1 is the first to extend this result to the setting where contextual node attributes are present.
+Notably, Braun et al.
+(2022) and Dreveton et al.
+(2024) also investigate contextual models similar to ours.
+However, neither Braun et al.
+(2022) nor Dreveton et al.
+(2024) allows edges to carry labels, and Dreveton et al.
+(2024) focuses specifically on achieving exact recovery.
+These can be viewed as special cases or reductions of our more general results (see Theorem 1 and Remark 1).
+In addition to the lower bound, we also design an efficient community detection algorithm for CLSBM.
+The algorithm begins by aggregating topological and contextual information into a latent factor model and then employs a spectral-based method to consistently estimate the community assignments for each node.
+1.1 Organization
+The remainder of the paper is organized as follows: In Section 2, we review related literature and highlight our main contributions.
+In Section 3.1, we formalize the model definitions and discuss the assumptions underpinning for our theoretical findings.
+Section 3.2 provides the theoretical lower bound for the number of misclassified nodes in the model and explores its connections with earlier work.
+Furthermore, Section 3.3 introduces an efficient spectral community detection algorithm for the proposed aggregated factor model and presents theoretical guarantees for its performance.
+Finally we conclude our results and future directions in Section 4.
+1.2 Notations
+In this paper, bold lowercase letters, such as $\mathbf{x}$, represent vectors, while bold uppercase letters, like $\mathbf{X}$, denote matrices and $3$-dimensional arrays.
 For a matrix $\mathbf{X}$, $\mathbf{X}_{ij}$ indicates the element located at the $i$-th row and $j$-th column, with $\mathbf{X}_{i\cdot}$ and $\mathbf{X}_{\cdot j}$ representing the $i$-th row and $j$-th column of $\mathbf{X}$, respectively.
-Similarly, for a $3$-d array $\mathbf{X}$, we use $\mathbf{X}_{i,j,k}$ to indicate the element at indices $i,j,k$ for each dimension.
+For a $3$-dimentional array $\mathbf{X}$, we use $\mathbf{X}(i,j,k)$ to indicate the element at indices $i,j,k$ for each dimension.
 Occasionally, $\mathbf{x}_{j}$ is used as a simplified notation for $\mathbf{X}_{\cdot j}$.
@@
 For any integer $K\in\mathbb{N}$, the set $\{1,2,3,\cdots,K\}$ is represented by $[K]$.
-The notation $\left\|\mathbf{x}\right\|_{q}$, $\left\|\mathbf{X}\right\|_{F}$, $\left\|\mathbf{X}\right\|_{op}$ represent the $\ell_{p}$ norm of a vector $\mathbf{x}$, Frobenius norm and operoator norm of matrix $\mathbf{X}$ respectively.
+The notation $\left\|\mathbf{x}\right\|_{q}$, $\left\|\mathbf{X}\right\|_{F}$, $\left\|\mathbf{X}\right\|_{op}$ represent the $\ell_{p}$ norm of a vector $\mathbf{x}$, Frobenius norm and operator norm of matrix $\mathbf{X}$ respectively.
 Here we use $\mathbf{1}(\cdot)$ to display the indicator function.
-For two sequences $a_{n}$ and $b_{n}$, the notation $a_{n}\lesssim b_{n}$ implies $a_{n}\leq Cb_{n}$ for some constant $C$.
-2 Related work and main contribution
-Jog and Loh (2015) was the first to investigate the threshold of exact recovery for the labeled/weighted SBM model by extending the results from Zhang and Zhou (2016).
-Additionally, Yun and Proutiere (2016) derived the optimal misclassification proportion for the LSBM and introduced a two-stage algorithm to achieve the information-theoretic limit.
-Besides, the contextual SBM, a generalization of SBM that incorporates additive node attributes to provide further information for identifying community clusteres, has been explored in various studies Braun et al.
+For two sequences $a_{n}$ and $b_{n}$, the notation $a_{n}\lesssim b_{n}$ means $a_{n}\leq Cb_{n}$ for some constant $C$.
+2 Related work and main contributions
+The threshold for exact recovery in the weighted/labeled SBM was first investigated by Jog and Loh (2015), which extends the results established in Zhang and Zhou (2016) for the SBM.
+Moreover, Yun and Proutiere (2016) derived the optimal misclassification rate for the LSBM and introduced a two-stage algorithm to achieve the information-theoretic limit.
+Besides, the Contextual-SBM, a generalization of SBM that incorporates additive node attributes to provide further information for identifying communities, has been explored in various studies Braun et al.
 (2022); Deshpande et al.
@@
 (2022).
-However, to the best of our knowledge, the optimal classification rate for models with general non-binary labeled edges combined with independent arbitrary node attributes still remains unexplored.
-Our Theorem 1 shows that the lower bound for the expected misclasification number of nodes (as defined in Eq.
+However, to the best of our knowledge, the optimal misclassification rate for models with general non-binary labeled edges combined with independent arbitrary node attributes still remains unexplored.
+Our Theorem 1 establishes that the lower bound for the expected number of misclassified nodes (as defined in Eq.
 2) is $n\exp\left(-(1+\epsilon_{n})D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})n\right)$ for some sequence $\epsilon_{n}$ approaching zero.
 Here the key quantity $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ (defined in (4)) is the divergence defined based on the minimal value for the sum of two KL divergences, representing the topological and attribute information respectively.
-Furthermore, Remark 1 reveals a connection between the minimal value of sum of KL divergences over constrained sets and the Chernoff divergence, extending the results of (Yun and Proutiere, 2016, Claim $4$) to our settings and consequently prove that the aformentioned lower bound indeed match the optimal classfication rate in (Label)SBM(Zhang and Zhou (2016); Yun and Proutiere (2016)) and Gaussian mixture model(Lu and Zhou (2016)).
-One of our main technical contribution stems in the analysis for the solve of the problem (4).
+Furthermore, Remark 1 reveals a connection between the minimal value of sum of KL divergences over constrained sets and the Chernoff divergence, extending the results of (Yun and Proutiere, 2016, Claim $4$) to our settings and consequently prove that the aformentioned lower bound indeed match the optimal misclassification rate in the LSBM (Zhang and Zhou (2016); Yun and Proutiere (2016)) and the Gaussian mixture model (Lu and Zhou (2016)).
+Our main technical contribution stems in the analysis to the problem (4).
 It is worth noting that, while the optimal solution for $D_{A}$ has been investigated in Yun and Proutiere (2016), the impact of the newly introduced term $D_{X}$ on the main results remains unclear.
-The additional term introduces further challenges, as it requires solving an optimization problem in function spaces, which is significantly more complex.
-In addition to establishing the theoretical lower bound, we also apply spectral methods to the proposed aggregated latent factor model, achieving a polynomial error rate.
-While this result is not optimal, the output of the spectral method can serve as a valuable initialization for further refinement in a second step, e.g see Gao et al.
+The additional term introduces further challenges, as it involves in solving an optimization problem in function spaces, which is significantly more complex.
+In addition to establishing the theoretical lower bound, we also apply spectral-based methods to the proposed aggregated latent factor model, achieving a polynomial misclassification rate.
+Despite the suboptimality of the result, the output of the spectral-based method provides a valuable initialization for further refinement in a subsequent step (e.g., Gao et al.
 (2018); Yun and Proutiere (2016); Braun et al.
-(2022).
+(2022)).
 3 Model description and main result
 3.1 Model definition
-We primarily follow the notation introduced in Yun and Proutiere (2016).
-Each node is independently assigned to a cluster $\mathcal{V}_{k}$ with probability $\alpha_{k}$, where $k\in[K]$, and this assignment is independent of the node number $n$.
-Each edge $A_{i,j}\in\mathcal{V}^{2}$ is labeled with $l\in\{0,1,\cdots,L\}$, where the labeling probability depends on $i$, $j$, and $l$, and is denoted as $\mathbf{P}(\sigma(i),\sigma(j),l)$.
-Here, $\mathbf{P}\in\mathbb{R}^{K\times K\times(L+1)}$, and $\sigma(\cdot)$ is the node cluster indicator function.
-Here we summerize the above in the following definition of LSBM model as follows:
-Definition 1 (Label Stochastic Block Model).
-Let $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$ be an $3$-d tensor and we write $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}\sim LSBM\left(L,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$ if $\forall l\in\{0,1,\cdots,L\}$ $\mathbf{A}_{l}\in\left\{0,1\right\}^{n\times n}$ is symmertical and its upper triangular has independent elements such that
-$\mathbb{P}(\mathbf{A}_{l}(i,j)=1)=\mathbf{P}(\sigma(i),\sigma(j),l),~{}\sum_{l=0}^{L}\mathbf{A}_{l}(i,j)=1,~{}\textit{and}\quad\mathbb{P}(\sigma(i)=k)=\boldsymbol{\alpha}_{k},$ (1)
-$\forall i\neq j\in[n],\,k\in[K]$.
-Here we assume $\mathbf{A}_{l}(i,i)=0$ for all $i\in[K],l\in[L]$ to remove the self loop of the graph, $\boldsymbol{\alpha}$ is a constant vector independent of $n$ and $\sigma(\cdot):[n]\rightarrow[K]$ is the cluster assignment indicator function.
-It is worth mentioning here that both $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$ and any third dimension of $\mathbf{P}$, such as $\mathbf{P}_{ij\cdot}\in\mathbb{R}_{+}^{(L+1)}$, must correspond to a probability distribution.
-In addition to topological information, contextual attributes are also incorporated into the final model.
-We assume that each node $i$ is equipped with an attribute vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$, and the network $\mathbf{A}$ is considered independent of the sets of attribute vectors $\{\mathbf{x}_{i}\}$ given the label function $\sigma(\cdot)$.
-This setting was initially proposed in Deshpande et al.
+We largely follow the notation introduced in Yun and Proutiere (2016) to describe the LSBM.
+Denote $\mathcal{V}$ as the set of nodes constrained to $\textit{card}(\mathcal{V})=n$ and let $\{\mathcal{V}_{k}\}_{k\in[K]}$ represent the $K$ disjoint communities.
+Here each node is independently assigned to a community $\mathcal{V}_{k}$ with probability $\alpha_{k}$, where $k\in[K]$ and $\mathbf{1}^{T}\mathbf{\alpha}=1$, and this assignment is independent of the number of nodes $n$.
+Each edge $(i,j)\in\mathcal{V}^{2}$ is assigned to a label $l\in\{0,1,\cdots,L\}$, where the probability of labeling depends on $i$, $j$, and $l$, and is denoted by $\mathbf{P}(\sigma(i),\sigma(j),l)$.
+Here, $\mathbf{P}\in\mathbb{R}^{K\times K\times(L+1)}$, and $\sigma(\cdot)$ is the ground truth indicator function.
+We summarize the above in the definition of LSBM as follows:
+Definition 1 (LSBM).
+Let $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$ and we denote $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}\sim\textit{LSBM}\left(L,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$ if $\forall l\in\{0,1,\cdots,L\}$ $\mathbf{A}_{l}\in\left\{0,1\right\}^{n\times n}$ is symmertical and its upper triangular elements are independent, such that:
+$\mathbb{P}(\mathbf{A}_{l}(i,j)=1)=\mathbf{P}(\sigma(i),\sigma(j),l),~{}\sum_{l=0}^{L}\mathbf{A}_{l}(i,j)=1,$
+$\mathbb{P}(\sigma(i)=k)=\alpha_{k},$ (1)
+$\forall i\neq j\in[n],\,k\in[K]$. Here we assume $\mathbf{A}_{0}(i,i)=1$ for all $i\in[K]$ to eliminate the self-loops within the graph.
+In addition to topological information, contextual node attributes are also incorporated into the final model.
+We assume that each node $i$ is associated with an attribute vector $\mathbf{x}_{i}\in\mathbb{R}^{d}$, and the graph $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}$ is considered independent of the sets of attribute vectors $\{\mathbf{x}_{i}\}_{i\in[n]}$ given the ground truth indicator function $\sigma(\cdot)$.
+This setting was initially proposed in Zhang et al.
 (2016) and has since been discussed in subsequent works Binkiewicz et al.
-(2017); Yan and Sarkar (2021); Weng and Feng (2016).
-Definition 2 (Contexual Label Stochastic Block Model).
-Let $d\geq 2$, $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$, $\boldsymbol{\mu}\in\mathbb{R}^{d\times K}$ ,and we write ($\left(\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}},\left\{\mathbf{x}_{i}\right\}_{i\in[n]}\right)\sim CLSBM(L,d,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot))$, if $\{\mathbf{A}_{l}\}_{l\in\{0,1,\cdots,L\}}\sim LSBM\left(n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$, and $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}\sim\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$.
-Here we assume $\left\{\mathbf{A}_{l}\right\}_{l\in[L]}$ and $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$ are independently generated given the parameters respectively.
-Assume there exist a assignment vector fucntion $\bar{\sigma}(\cdot)$ to indicate the assignment for each node and this $\bar{\sigma}(\cdot)$ is estimated from the observed labels $\left\{\mathbf{A}_{l}\right\}_{l\in[L]\bigcup\{0\}]}$ and attribute $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$.
-In the next section, we provide the information theoretical lower-bound for the misclassification rate between estimated $\bar{\sigma}(\cdot)$ and ground truth $\sigma(\cdot)$ such that
-$\bar{s}:=\inf_{\begin{subarray}{c}\pi\in S_{n}\\ \bar{\sigma}\in\mathcal{F}_{a}\end{subarray}}~{}\mathbb{E}\left[\sum_{i\in n}\mathbf{1}_{\{(\bar{\sigma}(\pi(i)))\neq\sigma(i))\}}\right].$ (2)
-Here, $S_{n}$ represents the set of all permutations of $[K]$, and $\mathcal{F}_{a}$ denotes the set of all assignment functions $\sigma:[N]\to[K]$.
-Before presenting the main results, we first outline our assumptions.
+(2017); Yan and Sarkar (2021); Weng and Feng (2016); Deshpande et al.
+(2018).
+Definition 2 (CLSBM).
+Let $d\geq 1$, $L\geq 1$, $K\geq 2$, $\boldsymbol{\alpha}\in\mathbb{R}_{+}^{K}$, $\mathbf{P}\in\mathbb{R}_{+}^{K\times K\times(L+1)}$, $\boldsymbol{\mu}\in\mathbb{R}^{d\times K}$, and we denote ($(\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}},\left\{\mathbf{x}_{i}\right\}_{i\in[n]})\sim\textit{CLSBM}(L,d,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot))$, if $\{\mathbf{A}_{l}\}_{l\in\{0,1,\cdots,L\}}\sim\textit{LSBM}\left(L,n,K,\mathbf{P},\boldsymbol{\alpha},\sigma(\cdot)\right)$, and $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}\sim\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$.
+Let $\bar{\sigma}(\cdot)$ be an assignment vector function that indicates the assignment for each node and is estimated from the observed edge labels $\{\mathbf{A}_{l}\}_{l\in\{0,1,\cdots,L\}}$ and node attributes $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$.
+In the next section, we derive an information-theoretic lower bound for the expected number of misclassified nodes, valid for all the possible choices of $\tilde{\sigma}(\cdot)$, defined as follows
+$\bar{s}:=\inf_{\begin{subarray}{c}\pi\in S_{K}\\ \tilde{\sigma}\in\mathcal{F}_{I}\end{subarray}}~{}\mathbb{E}\left[\sum_{i\in[n]}\mathbf{1}_{\{\pi(\tilde{\sigma}(i))\neq\sigma(i)\}}\right].$ (2)
+Here, $S_{K}$ represents the set of all permutations of $[K]$, and $\mathcal{F}_{I}$ denotes the set of all indicator functions.
+Before presenting the main results, we first outline our assumption as follows.
 Assumption 1.
 $\exists~{}\eta_{1},\eta_{2}>0$ such that
-$\forall l\in\{0,1,\cdots,L\},~{}\forall i,j,k\in[K],\quad\frac{\mathbf{P}(i,j,l)}{\mathbf{P}(i,k,l)}\leq\eta_{1},\quad\textit{and},\quad\left\|\boldsymbol{\mu}_{i}\right\|_{2}\leq\eta_{2}.$ (3)
-The assumption mentioned earlier arises from the limitations of our proof techniques, and we believe it could potentially be eliminated with improved proof mechanisms.
-Notably, the assumption regarding $\mathbf{P}$ originates from Yun and Proutiere (2016), while the assumption for $\boldsymbol{\mu}$ is introduced to bound the second-order moments of the pseudolikelihood estimator for the perturbed model in the proof of Theorem 1.
+$\forall l\in\{0,1,\cdots,L\},~{}\forall i,j,k\in[K],\quad\frac{\mathbf{P}(i,j,l)}{\mathbf{P}(i,k,l)}\leq\eta_{1},\quad\textit{and}\quad\left\|\boldsymbol{\mu}_{i}\right\|_{2}\leq\eta_{2}.$ (3)
+Although counterintuitive, the aforementioned assumption implies that the signal contributed by any one of the $K$ communities cannot be excessively strong for both the network and node attributes.
+This assumption arises from the limitations of our proof techniques, and it could potentially be eliminated with improved proof mechanisms.
+Specifically, the assumption for $\boldsymbol{\mu}$ is newly introduced here to bound the second-order moments of the pseudo-likelihood estimator for the perturbed model in the proof of Theorem 1.
+Similarly, the assumption regarding $\mathbf{P}$ originates from Yun and Proutiere (2016), serves the the same purpose.
 It is also worth mentioning that similar assumptions on $\boldsymbol{\mu}$ are commonly found in previous literature, such as Abbe et al.
 (2022).
-3.2 Lower bound on number of misclassfication nodes
-In this section we provide our main result. We start with the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$:
-$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}):=\min_{k_{1}\neq k_{1}\in[K]}\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}_{+}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})+\frac{1}{n}D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X})$ (4)
+3.2 Lower bound on the number of misclassification nodes
+In this section we provide one of our main results. We start with the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$:
+$D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}):=\min_{k_{1}\neq k_{2}\in[K]}\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}_{+}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})+\frac{1}{n}D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X}),$ (4)
 where
-$D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A}):=\max\left\{\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{1})\right),\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{2})\right)\right\},$
+$D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A}):=\max\biggl{\{}\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k)~{}\|~{}\mathbf{P}(k,k_{1})\right),\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k)~{}\|~{}\mathbf{P}(k,k_{2})\right)\biggr{\}},$
+$D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X}):=\max\biggl{\{}\operatorname{KL}_{X}\left(q_{X}~{}\|~{}p_{X}(k_{1})\right),\operatorname{KL}_{X}\left(q_{X}~{}\|~{}p_{X}(k_{2})\right)\biggr{\}}.$ (5)
 […]
-$D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X}):=\max\biggl{\{}\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{1})\right),\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{2})\right)\biggr{\}}.$ (5)
-[…]
-Theorem 1.
-Denote $\bar{p}:=\max_{i,j,l\geq 1}\mathbf{P}(i,j,l)$, grant $\bar{p}=\omega(1/n)$, $\bar{p}=o(1)$ and 1. One has
-lim $\liminf\frac{nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})}{\log(n/s)}\geq 1$ (6)
-if $\eta_{2}=o(n)$. Here we assume $\limsup_{n\rightarrow\infty}\frac{\bar{s}}{s}\leq 1$. Here $\bar{s}$ is defined in (2).
+Theorem 1 (Lower bound).
+Denote $\bar{p}:=\max_{i,j,l\geq 1}\mathbf{P}(i,j,l)$, grant 1 and assume $\bar{p}=\omega(1/n)$, $\bar{p}=o(1)$, and $\eta_{2}=o(n)$.
+Let $s=o(n)$.
+If there exists an algorithm that asymptotically has fewer misclassified nodes than $s$ in expectation, i.e., $\limsup_{n\rightarrow\infty}\frac{\bar{s}}{s}\leq 1$, then we have
+$\liminf_{n\rightarrow\infty}\frac{nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})}{\log(n/s)}\geq 1,$ (6)
+where $\bar{s}$ is defined in (2).
 Proof.
-See Appendix A. ∎
-Theorem 1 indeed implies that the number of misclassfied nodes for the proposed CLSBM model is above the order $n\exp(-nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}))$.
-The following remark indeed builds the connections between the Chernoff information $I$ proposed in (Dreveton et al., 2024, equation $3.4$) and $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ in (4)
+We defer the proof to Appendix A. ∎
+Theorem 1 implies that the expected number of misclassified nodes for the proposed CLSBM is at least of the order $n\exp(-nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}))$.
+The following remark indicates that the proposed measure $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ generalizes both the CH-divergence for the LSBM (Yun and Proutiere, 2016, Claim 4) and the minimal distance between the centers of Gaussian mixtures.
 Remark 1.
-When $\bar{p}=o(1)$, for C-LSBM model, one has
+When $\bar{p}=o(1)$, for CLSBM, one has
 $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$
-$=\min_{k_{1}\neq k_{2}}\max_{t\in[0,1]}\sum_{k}^{K}\alpha_{k}\biggl{[}\sum_{l=1}^{L}\left((1-t)\mathbf{P}(k,k_{1},l)+t\mathbf{P}(k,k_{2},l)-\mathbf{P}^{1-t}(k,k_{1},l)\mathbf{P}^{t}(k,k_{2},l)\right)+\frac{t(1-t)}{2n}\left\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}\biggr{]}.$ (7)
-Proof.
-See Appendix B. ∎
-From Remark 1, if the contextual attribute is missing such that $\boldsymbol{\mu}=0$, the divergence $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ reduces to the CH-divergence, as stated in Yun and Proutiere (2016).
-On the other hand, when the topological information is missing (e.g., $\mathbf{P}(k,k_{1},l)=\mathbf{P}(k,k_{2},l)$ for all $k,k_{1},k_{2},l$), we have $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})=\frac{1}{8n}\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\|_{2}^{2}$, which corresponds to the Gaussian mixture model, as shown in (Lu and Zhou, 2016, Theorem 3.3).
-3.3 Spectral clustering for latent factor model
-In this section, we extend the proposed CLSBM to a semiparametric model and demonstrate that the spectral method can consistently estimate the proposed estimator $s(\hat{\sigma})$ defined in (10), given the adjacency matrices $\{\mathbf{A}_{l}\}_{l\in\{0,1,2,\dots,L\}}$ and node attribute vectors $\{\mathbf{x}_{i}\}_{i\in[n]}$.
-Specifically, we first outline the construction of the estimator by combining the topological information $\{\mathbf{A}_{l}\}_{l\in\{0,1,2,\dots,L\}}$ with a Gaussian mixture model and provide an intuitive explanation of the spectral method at the population level in Section 3.3.1.
-Subsequently, in the following section, we formalize the spectral clustering method in Section 3.3.2 and establish the consistency of the proposed method, as proven in Lemma 1, to demonstrate its effectiveness.
+$\!\!=\min_{k_{1}\neq k_{2}\in[K]}\max_{t\in[0,1]}\sum_{k=1}^{K}\alpha_{k}\biggl{[}\sum_{l=1}^{L}\biggl{(}(1-t)\mathbf{P}(k,k_{1},l)+t\mathbf{P}(k,k_{2},l)-\mathbf{P}^{1-t}(k,k_{1},l)\mathbf{P}^{t}(k,k_{2},l)\biggr{)}\!\!+\frac{t(1-t)}{2n}\left\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}\biggr{]}.$ (7)
+The proof is deferred Appendix B.
+When the contextual node attributes are absent (i.e., $\boldsymbol{\mu}=0$), the divergence $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ simplifies to the CH-divergence introduced in Yun and Proutiere (2016).
+On the other hand, in cases where topological information is missing (e.g., $\mathbf{P}(k,k_{1},l)=\mathbf{P}(k,k_{2},l)$ for all $k,k_{1},k_{2},l$), we have $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})=\frac{1}{8n}\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\|_{2}^{2}$, which corresponds to threshold for the Gaussian mixture model, as shown in (Lu and Zhou, 2016, Theorem 3.3).
+3.3 Spectral community detection for latent factor model
+In this section, we extend the proposed CLSBM to a latent factor model and demonstrate that the spectral-based method can consistently estimate the proposed estimator $s(\hat{\sigma})$ defined in (10), based on the adjacency matrices $\{\mathbf{A}_{l}\}_{l\in\{0,1,2,\dots,L\}}$ and contextual node attribute vectors $\{\mathbf{x}_{i}\}_{i\in[n]}$.
+Specifically, we first outline the construction of the estimator by combining the topological information with the contextual node attributes and provide an intuitive explanation of the spectral-based method in Section 3.3.1.
+Subsequently, in the following section, we formalize the spectral-based method in Section 3.3.2 and establish its consistency, as proven in Lemma 1, thereby demonstrating its effectiveness.
 3.3.1 Aggregated latent factor model
-The latent factor model demonstrates its capability in revealing hidden low-dimensional structures within observed high-dimensional data and has been widely applied in various domains, including PCA, Gaussian mixtures, Nonnegative Matrix Factorization, and Sparse Factor Analysis.
-This provides an intuitive basis for combining both the topological model and Gaussian mixture into a shared latent space, as both the adjacency matrix and the observed Gaussian mixture share the same column space spanned by the ground truth assignment matrix $\mathbf{Z}(\sigma)\in\mathbb{R}^{n\times k}$, where $\mathbf{z}_{i\cdot}:=\mathbf{e}_{\sigma(i)}^{T}$.
-To elaborate, we begin by defining our aggregator factor matrix $\mathbf{S}$: given weights $\{w_{l}\}_{l\in[L]}$ independently and identically sampled from a uniform distribution over $[0,1]$, we define
-$\mathbf{S}:=\sum_{l=1}^{L}w_{l}\mathbf{A}^{(l)}+\frac{1}{n}\mathbf{X}\mathbf{X}^{T}.$ (8)
-where $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents the collection of $n$ observed Gaussian mixture data.
-The weights $\{w_{l}\}_{l\in[L]}$ are used to identify signals for different labels $l\in[L]$, and the factor $1/n$ balances the signal-to-noise ratio (SNR) between the graph signal and the Gaussian signal, as motivated by Remark 1.
-Let $\mathbf{P}_{s}:=\frac{1}{L}\sum_{l=1}^{L}\mathbf{P}_{\cdot,\cdot,l}\in\mathbb{R}^{K\times K}$.
-At the population level, we have:
+The latent factor model demonstrates its capability in revealing hidden low-dimensional structures within observed high-dimensional data and has been widely applied in various domains, including PCA, Gaussian mixtures, nonnegative matrix factorization, sparse factor analysis and so on.
+This provides an intuitive basis for combining both the topological model and contextual node attributes into a shared latent space, as both the adjacency matrix and the node attributes share the same eigenspace spanned by the ground truth assignment matrix $\mathbf{Z}(\sigma)\in\mathbb{R}^{n\times k}$, where $\mathbf{Z}_{i\cdot}:=\mathbf{e}_{\sigma(i)}^{T}$.
+To elaborate, we begin by defining our aggregator factor matrix $\mathbf{S}$: Given weights $\{w_{l}\}_{l\in[L]}$ i.i.d sampled from a uniform distribution over $[0,1]$, we define:
+$\mathbf{S}:=\sum_{l=1}^{L}w_{l}\mathbf{A}_{l}+\frac{1}{n}\mathbf{X}^{T}\mathbf{X},$ (8)
+where $\mathbf{X}\in\mathbb{R}^{d\times n}$ is a matrix whose columns represent the collection of $n$ observed node attributes.
+The weights $\{w_{l}\}_{l\in[L]}$ are employed to identify signals for different labels $l\in[L]$, and the factor $1/n$ balances the signal-to-noise ratio (SNR) between the graph signal and the Gaussian signal, as motivated by Remark 1.
+Let $\mathbf{P}_{s}:=1/(2L)\sum_{l=1}^{L}\mathbf{P}(\cdot,\cdot,l)\in\mathbb{R}^{K\times K}$.
+At the population level, the following holds:
 $\mathbb{E}\left[\mathbf{S}\right]$ $=\mathcal{H}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})+\frac{1}{n}\mathbf{Z}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\mathbf{Z}^{T}$
-$=\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}-\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})+\frac{1}{n}\mathbf{Z}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\mathbf{Z}^{T}$
-$=\underbrace{\mathbf{Z}\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)\mathbf{Z}^{T}}_{\mathbf{M}}-\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}).$ (9)
-Here, $\mathcal{H}(\cdot)$ denotes the hollowing operation, which removes the diagonal elements of the matrix, and $\mathbf{Z}$ is shorthand for $\mathbf{Z}(\sigma)$.
-It is evident that the first term $\mathbf{M}$ in $\mathbb{E}\left[\mathbf{S}\right]$ is a symmetric low-rank factor model, which has been extensively studied in areas such as low-rank matrix decomposition, Factor Analysis, signal processing, and spectral clustering.
-By the definition of $\mathbf{M}$, it is clear that the optimal solution of the $k$-means algorithm applied to the columns/rows of $\mathbf{M}$ should recover the ground truth assignment $\mathbf{Z}$.
-In practice, however, instead of directly applying $k$-means to the columns of $\mathbf{M}$ (which has dimension $n$), spectral methods are often employed due to their computational advantages.
-Specifically, let $\mathbf{U}_{K}$ denote the top $K$ eigenvectors of $\mathbf{M}$.
+$=\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}-\operatorname{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})+\frac{1}{n}\mathbf{Z}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\mathbf{Z}^{T}$
+$=\underbrace{\mathbf{Z}\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)\mathbf{Z}^{T}}_{\mathbf{M}}-\operatorname{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T}).$ (9)
+Here, $\mathcal{H}(\cdot)$ denotes the hollowing operation, which sets the diagonal elements of the matrix to zero, and $\mathbf{Z}$ is shorthand for $\mathbf{Z}(\sigma)$.
+It is evident that the first term $\mathbf{M}$ in $\mathbb{E}\left[\mathbf{S}\right]$ is a symmetric low-rank factor model when $K<n$, which has been extensively studied in areas such as low-rank matrix decomposition, factor analysis, signal processing, and spectral clustering.
+By definition, the optimal solution of the $k$-means algorithm applied to the columns/rows of $\mathbf{M}$ should yeild the ground truth assignment $\mathbf{Z}$.
+In practice, however, instead of directly applying $k$-means to the columns of $\mathbf{M}$ (which has dimension $n$), spectral methods Newman (2006) are often pre-employed due to their computational effficiency.
+Specifically, let $\mathbf{U}_{K}\in\mathbb{R}^{n\times K}$ denote the top-$K$ eigenvectors of $\mathbf{M}$.
 We apply $k$-means to the columns of $\mathbf{U}_{K}^{T}\mathbf{M}$.
-This is equivalent to applying $k$-means to the columns of $\mathbf{M}$ or $\mathbf{U}_{K}^{T}\mathbf{M}$ in terms of clustering outcomes, as $\mathbf{U}_{K}$ is an orthonormal matrix that preserves distances during projection.
-However, since neither $\mathbf{M}$ nor $\mathbf{U}_{K}$ is observed, even at the population level, we substitute $\mathbf{S}$ for $\mathbf{M}$ and use its top eigenvectors, denoted as $\bar{\mathbf{U}_{K}}$, in place of $\mathbf{U}_{K}$.
-Given that the second term in (3.3.1) is small in operator norm, i.e., $\|\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})\|_{\text{op}}=o(1)$, the misclassification rate of the $k$-means algorithm based on $\mathbf{S}$ depends on $\|\mathbf{S}-\mathbb{E}(\mathbf{S})\|_{F}$, which can be bounded using concentration inequalities.
-We summarize the spectral clustering procedure in Algorithm 1.
-1:Node attributes $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$, adjacent matrix $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}$, and number of clusters $K$
-2:Cluster assignments for each data point
+In fact applying $k$-means to the columns of $\mathbf{M}$ or $\mathbf{U}_{K}^{T}\mathbf{M}$ yeilds equivalent clustering outcomes, as $\mathbf{U}_{K}$ is an orthonormal matrix that preserves distances during projection.
+Chooseing $\mathbf{U}_{K}^{T}\mathbf{M}$ is motivated by its computational efficiency for clustering, particularly because $K$ is typically much smaller than $n$.
+However, since neither $\mathbf{M}$ nor $\mathbf{U}_{K}$ is observed even at the population level, we substitute $\mathbf{S}$ for $\mathbf{M}$ and employ its top-$K$ eigenvectors, denoted as $\bar{\mathbf{U}}_{K}$, in place of $\mathbf{U}_{K}$.
+Given that the second term in (3.3.1) has a small operator norm, i.e., $\|\text{diag}(\mathbf{Z}\mathbf{P}_{s}\mathbf{Z}^{T})\|_{\text{op}}=o(1)$, the misclassification rate of the $k$-means algorithm based on $\mathbf{S}$ depends on $\|\mathbf{S}-\mathbb{E}(\mathbf{S})\|_{F}$, a quantity can be further bounded via some concentration inequalities.
+We summarize the aforementioned procedures in Algorithm 1, refer to as Spectral Community Detection.
+1:Node attributes $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$, adjacency matrices $\left\{\mathbf{A}_{l}\right\}_{l\in\{0,1,\cdots,L\}}$, and number of communities $K$.
+2:Assignment for each node
 3:Step 1: Compute eigenvectors
@@
 6:Denote $\mathbf{Q}:=\bar{\mathbf{U}}_{K}^{T}\mathbf{S}\in\mathbb{R}^{k\times n}$ and denote $\bar{\sigma}(\cdot)$ as the solution to $k$-means such that:
-$\bar{\sigma}(\cdot),\left\{\bar{\boldsymbol{\theta}}\right\}_{i\in[K]}=\operatorname*{argmin}_{\begin{subarray}{c}\sigma\in\mathcal{F}_{a},\\ \boldsymbol{\theta}_{i}\in\mathbb{R}^{k}\end{subarray}}\sum_{i\in[n]}\left\|\mathbf{Q}_{i}-\boldsymbol{\theta}_{\sigma(i)}\right\|_{2}^{2}.$
-7:Return Cluster assignments function $\bar{\sigma}(\cdot)$.
-Algorithm 1 Spectral Clustering Algorithm
+$\bar{\sigma}(\cdot),\left\{\bar{\boldsymbol{\theta}}\right\}_{i\in[K]}=\operatorname*{argmin}_{\begin{subarray}{c}\sigma\in\mathcal{F}_{a},\\ \boldsymbol{\theta}_{i}\in\mathbb{R}^{k}\end{subarray}}\sum_{i\in[n]}\left\|\mathbf{Q}_{\cdot i}-\boldsymbol{\theta}_{\sigma(i)}\right\|_{2}^{2}.$
+7:Return: Assignment indicator function $\bar{\sigma}(\cdot)$.
+Algorithm 1 Spectral Community Detection
 3.3.2 Theoretical guarantee
-Denote $s(\bar{\sigma})$ as the number of misclassified nodes for label function $\bar{\sigma}$ recovered from Algorithm 1 such that
-$s(\bar{\sigma}):=\inf_{\pi\in S_{n}}~{}\mathbb{E}\left[\sum_{i\in n}\mathbf{1}_{\{(\bar{\sigma}(\pi(i)))\neq\sigma(i))\}}\right].$ (10)
+Denote $s(\bar{\sigma})$ as the number of misclassified nodes for indicator function $\bar{\sigma}$ recovered from Algorithm 1 such that
+$s(\bar{\sigma}):=\inf_{\pi\in S_{K}}~{}\mathbb{E}\left[\sum_{i\in n}\mathbf{1}_{\{\pi(\bar{\sigma}(i))\neq\sigma(i))\}}\right].$ (10)
 The following lemma thus provides the upper bound of $s(\bar{\sigma})$.
 Lemma 1.
-Grant the assumptions in theorem 1 and further assume $\bar{p}=w(\log(n)/n)$, let $\bar{\sigma}$ be the solution of the spectral algorithm, we have
-$s(\bar{\sigma})\lesssim\frac{K}{n\textit{SNR}}$ (11)
-with high probability.
-Here $\textit{SNR}:=(1+o(1))\min_{k_{1}\neq k_{2}}\left\|\left(\mathbf{P}_{s}\right)_{k_{1}}-\left(\mathbf{P}_{s}\right)_{k_{2}}+\frac{1}{n}\left(\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right)^{T}\boldsymbol{\mu}\right\|_{2}^{2}$.
-Furthermore if $n\textit{SNR}=w(1)$ we obtain $s(\bar{\sigma})=o(1)$.
+Grant the assumptions in Theorem 1 and further assume $\bar{p}=\omega(\log(n)/n)$, let $\bar{\sigma}$ be the solution of the Algorithm 1, we have
+$s(\bar{\sigma})\lesssim\frac{K}{n\cdot\mathrm{SNR}}$ (11)
+with high probability, where $\mathrm{SNR}:=\min_{k_{1}\neq k_{2}}\left\|\left(\mathbf{P}_{s}\right)_{k_{1}}-\left(\mathbf{P}_{s}\right)_{k_{2}}+\frac{1}{n}\left(\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right)^{T}\boldsymbol{\mu}\right\|_{2}^{2}$.
+Furthermore if $n\cdot\mathrm{SNR}=w(1)$ we obtain $s(\bar{\sigma})=o(1)$.
 Proof.
-See Appendix C
-Lemma 10 demonstrates that the proposed indicator $s(\bar{\sigma})$ is upper bounded by the inverse of the defined (SNR), which corresponds to the minimum distance between any two columns in the matrix $\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}$.
+The proof is deferred to Appendix C. ∎
+Lemma 1 demonstrates that the proposed indicator $s(\bar{\sigma})$ is upper bounded by the inverse of the defined (SNR) in a polynomial manner, which corresponds to the minimum distance between any two columns in the matrix $\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}$.
 Rates of this nature have been extensively studied in the literature Lei and Rinaldo (2015); Löffler et al.
-(2021); Zhang and Zhou (2024) for spectral methods under SBM or Gaussian mixtures, respectively.
-However, our work is the first to provide theoretical guarantees for spectral methods under our proposed CLSBM.
-Despite achieving consistency, our analysis does not attain the optimal rate presented in Theorem 1, which is exponentially small.
-However, we beileive that this estimate could indeed play the role of a initializers for the potential refinement as seen in Yun and Proutiere (2016); Gao et al.
+(2021); Zhang and Zhou (2024) for spectral-based methods under SBM or Gaussian mixtures, respectively.
+However, our work is the first to provide theoretical guarantees for spectral-based methods under the CLSBM.
+Despite achieving consistency, our analysis does not lead to the optimal rate presented in Theorem 1, which is exponentially small.
+However, we believe that this estimate can serve as an effective initializer for potential subsequent refinement steps, as demonstrated in Yun and Proutiere (2016); Gao et al.
 (2018).
+By leveraging the spectral method’s output as a starting point, further refinement techniques such as MLE can be applied to improve the accuracy of community detection, potentially bridging the gap between polynomial and exponential misclassification rates.
 4 Conclusion
@@
 Appendix A Proof of Theorem 1
-Mimic the change of measure method in Yun and Proutiere (2016) we adopt it to the CLSBM model and its pertubted version in this section.
-Denote $\Phi$ and $\Psi$ as the parameter for the true parameter model(CLSBM) and one potential perturbed parameter model respectively.
-To be more specific, the true ground truth model $\Phi$ first generate the cluster for each node based on $\boldsymbol{\alpha}$ such that $\mathbb{P}(\sigma(i)=k)=\alpha_{k}$ for the $i$-th node indepedently.
-For the topological imformation, model $\Phi$ generate the observed label $l\in[L]$ for two different node $i$, $j$ based on the $\mathbf{P}(\sigma(i),\sigma(j),l)$.
-For the contexual information, $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$.
+Mimicking the change of measure method in Yun and Proutiere (2016) we adapt it to the CLSBM model and its perturbed version in this section.
+Denote $\Phi$ and $\Psi$ as the parameter for the true parameter model (CLSBM) and a potential perturbed parameter model respectively.
+To be more specific, the true ground truth model $\Phi$ first generates the cluster for each node based on $\boldsymbol{\alpha}$ such that $\mathbb{P}(\sigma(i)=k)=\alpha_{k}$ for the $i$-th node indepedently.
+For the topological information, model $\Phi$ generates the observed labels $l\in[L]$ for two different nodes $i$, $j$ based on the $\mathbf{P}(\sigma(i),\sigma(j),l)$.
+For the contextual information, $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$.
 We state the definition of $\Psi$ as follows:
@@
 $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})=D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star})+\frac{1}{n}D_{X}(k_{1}^{\star},k_{2}^{\star},q_{X}^{\star})$ (12)
-The edge label generation mechanism for the topological information($\mathbf{A}$) in $\Psi$ is the same as in Yun and Proutiere (2016) by replace $\mathbf{q}$ with $\mathbf{q}_{\mathbf{A}}^{\star}$, here we state the generation of label for side information $\{\mathbf{x}_{i}\}_{i\in[n]}$ as follows: denote $i^{\star}=\operatorname*{argmin}_{\sigma(i)\in\{k_{1}^{\star},k_{2}^{\star}\}}i$.
-  • When $i\neq i^{\star}$: $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$ which is same as in $\Phi$.
+The edge label generation mechanism for the topological information($\mathbf{A}$) in $\Psi$ is the same as in Yun and Proutiere (2016) with replace $\mathbf{q}$ with $\mathbf{q}_{\mathbf{A}}^{\star}$, Here, we describe the generation of labels for the side information $\{\mathbf{x}_{i}\}_{i\in[n]}$ as follows: let $i^{\star}=\operatorname*{argmin}_{\sigma(i)\in\{k_{1}^{\star},k_{2}^{\star}\}}i$.
+  • When $i\neq i^{\star}$: $\mathbf{x}_{i}$ is sampled from $\mathcal{N}(\boldsymbol{\mu}_{\sigma(i)},\mathbf{I}_{d})$ which is the same as in $\Phi$.
   • When $i=i^{\star}$: $\mathbf{x}_{i^{\star}}$ is sampled from the distribution whose pdf is $q_{\mathbf{X}}^{\star}$.
-We denote the $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the pseudolikelihood ratio of the observed labels and contexual variables such that
+We denote $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the pseudolikelihood ratio of the observed labels and contextual variables such that:
 $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}:=\sum_{i\neq i^{\star}}^{n}\log\frac{\mathbf{q}_{A}^{\star}(\sigma(i),\mathbf{A}_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}.$ (13)
-In the following two sections we thus provide the detailed proof of theorem 1, which follows the proof in Yun and Proutiere (2016).
-The proof consist of two parts, the first part constructs a stochastic model $\Psi$ and lower bound mis-classification node $\mathbb{E}_{\Phi}(\epsilon^{\pi}(n))$ by the expectation of the proposed log-likelihood ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ defined in (13); In the second part we further estimate the ratio upper bound of $\mathbb{E}_{\Psi}(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}})$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]\right)^{2}\right]$ respectively to finish the proof.
+In the following two sections, we provide the detailed proof of theorem 1, which follows the proof in Yun and Proutiere (2016).
+The proof consists of two parts, the first part constructs a stochastic model $\Psi$ and lower bounds the misclassification error $\mathbb{E}_{\Phi}(\epsilon^{\pi}(n))$ using the expectation of the proposed log-likelihood ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ defined in (13); In the second part, we further estimate upper bounds for $\mathbb{E}_{\Psi}(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}})$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]\right)^{2}\right]$ respectively, to complete the proof.
 A.1 Part 1
-Leverage Lemma 2 we assert there exist $(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star},q_{X}^{\star})$ such that
+Leverage Lemma 2 we assert there exist $(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star},q_{X}^{\star})$ such that:
 $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ $=\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{1}^{\star})\right)+\frac{1}{n}\operatorname{KL}_{X}\left(q_{X}^{\star},p_{X}(k_{1}^{\star})\right)$
 $=\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}^{\star}(k),\mathbf{P}(k,k_{2}^{\star})\right)+\frac{1}{n}\operatorname{KL}_{X}\left(q_{X}^{\star},p_{X}(k_{2}^{\star})\right)$ (14)
-Recall the definition of the log ratio Eq. 13:
+Recall the definition of the log ratio Eq. 13 we obtain:
 $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}:=\sum_{i\neq i^{\star}}^{n}\log\frac{\mathbf{q}_{A}^{\star}(\sigma(i),\mathbf{A}_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}.$ (15)
-Mimic the approach described in Yun and Proutiere (2016), we could reach inequality $(17)$ in Yun and Proutiere (2016) adopted with our new defined log-ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the necessary condition for $\mathbb{E}[\epsilon(n)]\leq s$:
+Mimicking the approach described in Yun and Proutiere (2016), we can derive inequality $(17)$ in Yun and Proutiere (2016) adapted with our new defined log-ratio $\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ as the necessary condition for $\mathbb{E}[\epsilon(n)]\leq s$:
 $\log(n/s)-\log(2/\alpha_{k_{1}})\leq\mathbb{E}_{\Psi}(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}})+\sqrt{\frac{4}{\alpha_{k_{1}}}\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]\right)^{2}\right]}.$ (16)
-Then the next part finishes the proof by estiamting the bounds of $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}\right)^{2}\right]$ respectively.
+Then the next part finishes the proof by estimating the bounds of $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$ and $\mathbb{E}_{\Psi}\left[\left(\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}-\mathbb{E}_{\Psi}\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}\right)^{2}\right]$ respectively.
 A.2 Part 2
-  • Upper-bound for $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$:
+  • Upper bound for $\mathbb{E}_{\Psi}[\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}]$:
     Denote $L:=\log{\frac{\mathrm{d}\mathbb{P}_{\Psi}}{\mathrm{d}\mathbb{P}_{\Phi}}}$ and we obtain
@@
     $=\underbrace{\mathbb{E}_{\Psi}\sum_{i\neq i^{\star}}^{n}\log\frac{q_{A}^{\star}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}}_{\Gamma_{1}}+\underbrace{\mathbb{E}_{\Psi}\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}}_{\Gamma_{2}}$ (17)
-    $\Gamma_{1}$ has been carefully studied in (Yun and Proutiere, 2016, Theorem 1) such that
+    $\Gamma_{1}$ has been carefully studied in (Yun and Proutiere, 2016, Theorem 1) such that:
     $\Gamma_{1}\leq(n+2\log^{2}(n)\log(\eta_{1}))D_{A}(k_{1}^{\star},k_{2}^{\star},\mathbf{q}_{A}^{\star}).$
@@
     $\mathbb{E}_{\Psi}[L]\leq(n+2\log^{2}(n)\log(\eta_{1}))D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}).$
-    Furthermore repeating the procedure in (A.2) leads to
+    Furthermore, repeating the procedure in (A.2) leads to
     $\mathbb{E}_{\Psi}[L]\geq(n-2\log^{2}(n)\log(\eta_{1}))D(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu}).$
@@
     $\leq\mathbb{E}_{\Psi}[L^{2}]-\left(n-2\log^{2}(n)\right)^{2}D^{2}(\boldsymbol{\alpha},\mathbf{p},\mu).$ (19)
-    Then the following part remains to upper bound $\mathbb{E}_{\Psi}[L^{2}]$: Without loss of generaloity we assume $\sigma(i^{\star})=k_{1}$, by definition we have
+    Thus, the following part remains to upper bound $\mathbb{E}_{\Psi}[L^{2}]$: Without loss of generality we assume $\sigma(i^{\star})=k_{1}$, by definition we have
     $\mathbb{E}_{\Psi}[L^{2}]$
@@
     $\lesssim\frac{\log^{2}(\eta_{1})}{n^{2}}+\frac{\eta_{2}^{4}+\eta_{2}^{2}}{n^{4}}.$ (21)
-    Here we use $\mathbb{P}(i^{\star}>\log^{2}(n))\leq n^{-4}$ in the second step, independence of $\{\mathbf{x}\}_{i\in[L]}$ and $\mathbf{A}$ in the third step, equation (33) and assumption 1 in the fourth steps and $10<t<$ 1 in the last step.
+    Here, we use $\mathbb{P}(i^{\star}>\log^{2}(n))\leq n^{-4}$ in the second step, the independence of $\{\mathbf{x}\}_{i\in[L]}$ and $\mathbf{A}$ in the third step, equation (33) and 1 in the fourth step and $10<t<$ 1 in the last step.
 To proceeds we then focus on $\Gamma_{1}$:
@@
     $\quad+\left(\log(n)\log(\eta_{1})+\eta_{2}\right)^{2}+2\left(\log(n)\log(\eta_{1})+\eta_{2}\right)D(\boldsymbol{\alpha},\mathbf{p},\mu).$ (22)
-    Here we use the fact that any two nodes $i_{1}$ and $i_{2}$ such that $i_{1}\neq i_{2}\geq i^{\star}$ are muturally independent in the derivation. To upper bound $\Gamma_{11}$ we first have
+    Here we use the fact that any two nodes $i_{1}$ and $i_{2}$ satisfying $i_{1}\neq i_{2}\geq i^{\star}$ are mutually independent in the derivation. To upper bound $\Gamma_{11}$, we first have
     $\left(\sum_{i>i^{\star}}^{n}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}$
@@
     $\quad+2\left(\frac{n-i^{\star}}{n}\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star})}\right)\sum_{i>i^{\star}}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}$ (23)
-    We thus finish upper-bounding $\Gamma_{1}$ by analyzing $\Gamma_{11}$:
+    We thus complete the upper bound of $\Gamma_{1}$ by analyzing $\Gamma_{11}$:
     $\Gamma_{11}$ $=\mathbb{E}_{\Psi}\biggl{[}\left(\sum_{i>i^{\star}}^{n}\log\frac{q_{A}(\sigma(i),A_{i,i^{\star}})}{\mathbf{P}(\sigma(i),\sigma(i^{\star}),A_{i,i^{\star}})}+\log\frac{q_{X}^{\star}}{p_{X}(\sigma(i^{\star}))}\right)^{2}\Bigg{|}~{}\sigma(i^{\star})=k_{1},i^{\star}\leq\log^{2}(n)\biggr{]}$
@@
     $\lesssim n^{2}D^{2}(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu})+\left(n\log(\eta_{1})+\eta_{2}\right)D(\boldsymbol{\alpha},\mathbf{p},\boldsymbol{\mu})$ (25)
-    Further we end the proof by plugging above inequation back to (A.2):
+    Further we end the proof of upper bounding by plugging above inequalities back to (A.2):
     $\mathbb{E}_{\Psi}\left[\left(\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}-\mathbb{E}_{\Psi}[\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}]\right)^{2}\right]$
@@
 lim $\liminf\frac{nD(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})}{\log(n/s)}\geq 1.$
-Our proof then ends here.
+Our proof concludes here.
 Appendix B Proof of Remark 1
@@
 Proof.
-Recall the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$, pick any $k_{1}\neq k_{2}\in[K]$ and denote $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ such that
+Recall the definition of $D(\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$.
+For any $k_{1}\neq k_{2}\in[K]$ define $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ such that:
 $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu}):=\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})+\frac{1}{n}D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X}).$ (27)
-Here $D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})$ and $D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X})$ follow the definition in (3.2). Thus the proof remains to display
+Here, $D_{A}(k_{1},k_{2},\mathbf{P},\mathbf{q}_{A})$ and $D_{X}(k_{1},k_{2},\boldsymbol{\mu},q_{X})$ follow the definition in (3.2). Thus the proof reduces to showing:
 $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$
 $=\max_{t\in[0,1]}\sum_{k}^{K}\alpha_{k}\biggl{[}\sum_{l=1}^{L}\left((1-t)\mathbf{P}(k,k_{1},l)+t\mathbf{P}(k,k_{2},l)-\mathbf{P}^{1-t}(k,k_{1},l)\mathbf{P}^{t}(k,k_{2},l)\right)+\frac{t(1-t)}{2n}\left\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}\biggr{]}.$ (28)
-We start the proof by studying the cosntrained version of (27): Notice that $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ actually reach the minimum of the following problem:
+We begin the proof by studying the constrained version of (27): Note that $D^{\prime}(k_{1},k_{2},\boldsymbol{\alpha},\mathbf{P},\boldsymbol{\mu})$ actually reaches the minimum of the following problem:
 $\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{1})\right)+\frac{1}{n}\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{1})\right)$
@@
 $\quad>\sum_{k=1}^{K}\alpha_{k}\operatorname{KL}_{A}\left(\mathbf{q}_{A}(k),\mathbf{P}(k,k_{2})\right)+\frac{1}{n}\operatorname{KL}_{X}\left(q_{X},p_{X}(k_{2})\right).$ (29)
-Thus we write the lagrange dual form of Appendix B such that
+Thus, we express the lagrange dual form of Appendix B such that:
 $\max_{t\geq 0}\min_{\begin{subarray}{c}\mathbf{q}_{A}\in\mathbb{R}^{K\times L}\\ q_{X}\in\mathcal{F}\end{subarray}}L(\mathbf{q}_{A},q_{X},t)$ (30)
@@
 for some postive variable $t>0$.
-Notice that $\Gamma_{1}\left(\mathbf{q}_{A},t\right)$ has been completely studied in claim $4$ in Yun and Proutiere (2016), due to the independence of the $\mathbf{q}_{A},q_{X}$, the remaining part then focues on solving $\Gamma_{2}(q_{X},t)$.
-Since $q_{X}$ represent a probability density function that lies in the function space $\mathcal{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}$.
-Denote $\frac{\delta F(p)}{\delta p}(\mathbf{x})$ as the functional derivative of $F(p)$ to $p$, $\forall p\in\mathcal{F}$, by definition we have
+Notice that $\Gamma_{1}\left(\mathbf{q}_{A},t\right)$ has been completely studied in Claim $4$ in Yun and Proutiere (2016), due to the independence of the $\mathbf{q}_{A},q_{X}$, the remaining part then focuses on solving $\Gamma_{2}(q_{X},t)$.
+Since $q_{X}$ represents a PDF that lies in the function space $\mathcal{F}:\mathbb{R}^{d}\rightarrow\mathbb{R}$.
+We denote $\frac{\delta F(p)}{\delta p}(\mathbf{x})$ as the functional derivative of $F(p)$ with respect to $p$, $\forall p\in\mathcal{F}$.
+By definition, we have
 $\frac{\delta\Gamma_{2}(q_{X},t)}{\delta q_{X}}(\mathbf{x})$
@@
 Here we use $\frac{\delta\operatorname{KL}(q,p)}{\delta q}(\mathbf{x})=\log(q(\mathbf{x}))+1-\log(p(\mathbf{x}))$ in the derivation.
-Denote $q_{X}^{\star}(x):=\operatorname*{argmin}_{q_{X}\in\mathcal{F}}\Gamma_{2}(q_{X},t)$ and let (B) equal to zero we have:
+Let $q_{X}^{\star}(x):=\operatorname*{argmin}_{q_{X}\in\mathcal{F}}\Gamma_{2}(q_{X},t)$.
+Setting (B) equal to zero, we have:
 $\log(q_{X}^{\star}(\mathbf{x}))=(1-t)\log(p_{X}(k_{1})(\mathbf{x}))+t\log(p_{X}(k_{2})(\mathbf{x}))-1.$
 due to the convexity of $\Gamma_{2}(q_{X},t)$.
-Notice that $p_{X}(k_{1})$ and $p_{X}(k_{2})$ represent the pdf for $\mathcal{N}(\boldsymbol{\mu}_{k_{1}},\mathbf{I}_{d})$ and $\mathcal{N}(\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ respectively, consequently we obtain
+Note that $p_{X}(k_{1})$ and $p_{X}(k_{2})$ represent the PDF for $\mathcal{N}(\boldsymbol{\mu}_{k_{1}},\mathbf{I}_{d})$ and $\mathcal{N}(\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ respectively.
+Consequently, after some calculation, we obtain
 $q_{X}^{\star}(\mathbf{x})=C\exp(-\left\|\mathbf{x}-(1-t)\boldsymbol{\mu}_{k_{1}}-t\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}/2).$ (33)
-after some calculation. Here $C$ is some constant that is irelevent with $\mathbf{x}$. We assert $C=(2\pi)^{-d/2}$ by utilzing $q_{X}$ is an probability density function.
-Plug the expression of $q_{X}^{\star}$ back to $\Gamma_{2}(q_{X},t)$ defined in (B) gives:
+after some calculation. Here $C$ is a constant that is irelevent with $\mathbf{x}$. We assert $C=(2\pi)^{-d/2}$ by utilizing $q_{X}$ is an probability density function.
+Substituting the expression of $q_{X}^{\star}$ into $\Gamma_{2}(q_{X},t)$ defined in (B) yields:
 $\Gamma_{2}(q_{X}^{\star},t)=\frac{t(1-t)}{2}\left\|\boldsymbol{\mu}_{k_{1}}-\boldsymbol{\mu}_{k_{2}}\right\|_{2}^{2}.$ (34)
 after some algebra.
-Here we use Lemma 3 and the fact that $q_{X}^{\star}$ is the PDF for $\mathcal{N}((1-t)\boldsymbol{\mu}_{k_{1}}+t\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ in the derivation.
-We thus finish the proof by combining the expression of $\Gamma_{1}(\mathbf{q}_{A},t)$(see (Yun and Proutiere, 2016, Claim $4$)) and $\Gamma_{2}(q_{X}^{\star},t)$ together.
+Here, we use Lemma 3 and the fact that $q_{X}^{\star}$ is the PDF for $\mathcal{N}((1-t)\boldsymbol{\mu}_{k_{1}}+t\boldsymbol{\mu}_{k_{2}},\mathbf{I}_{d})$ in the derivation.
+We thus finish the proof by combining the expressions of $\Gamma_{1}(\mathbf{q}_{A},t)$(see (Yun and Proutiere, 2016, Claim $4$)) and $\Gamma_{2}(q_{X}^{\star},t)$ together.
 Appendix C Proof of Lemma 1
@@
 with probability at least $1-cn^{-c^{\prime}}$.
-Here we use $\mathbf{S}^{(k)}$ is the best approximation of $\mathbf{S}$ in the third step, (Coja-Oghlan, 2010, Lemma 8.5) and (Löffler et al., 2021, Lemma B.1) in the last step.
+Here we use tha fact that $\mathbf{S}^{(K)}$ is the best approximation of $\mathbf{S}$ in the third step, (Coja-Oghlan, 2010, Lemma 8.5) and (Löffler et al., 2021, Lemma B.1) in the last step.
 Notice that $\operatorname{rank}~{}(\mathbf{S}^{(K)}-\mathbf{M})\leq\operatorname{rank}(\mathbf{S}^{(K)})+\operatorname{rank}(\mathbf{M})\leq 2K$ we further have
@@
 $=\min_{\sigma(i)\neq\sigma(j)}\left\|\boldsymbol{\Sigma}\left(\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)_{\sigma(i)}-\left(\mathbf{P}_{s}+\frac{1}{n}\boldsymbol{\mu}^{T}\boldsymbol{\mu}\right)_{\sigma(j)}\right)\right\|_{2}$ (42)
-where $\boldsymbol{\Sigma}:=\operatorname{diag}([n_{1},n_{2},\cdots,n_{K}])=n\left(\operatorname{diag}(\boldsymbol{\alpha})+o(1)\mathbf{I}_{K}\right)$.
+where $\boldsymbol{\Sigma}:=\operatorname{diag}([n_{1},n_{2},\cdots,n_{K}])=n\left(\operatorname{diag}(\boldsymbol{\alpha})+\mathbf{\Delta})\mathbf{I}_{K}\right)$ for some $\mathbf{\Delta}$ i.e.
+$\left\|\mathbf{\Delta}\right\|_{op}=o(1)$.
 Here we use the orthonormality for the columns of $\mathbf{Z}\boldsymbol{\Sigma}^{-1}$ in the last step.
@@
 Lemma 3 (KL divergence for two Gaussian distributions).
-Consider two Gaussian probability distributions with unit variance, having means $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$, and probability density functions (pdfs) denoted as $p_{1}$ and $p_{2}$, respectively.
+Consider two Gaussian probability distributions with unit variance, having means $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$, and PDFs denoted as $p_{1}$ and $p_{2}$, respectively.
 It follows that one has
@@
 $\hat{\sigma}(\cdot),\{\mathbf{c}_{i}\}_{i\in[n]}:=\operatorname*{argmin}_{\begin{subarray}{c}\hat{\sigma}(\cdot)\in\mathcal{F},\\ \mathbf{c}_{i}\in\mathbb{R}^{n}\end{subarray}}\sum_{i\in[n]}\left\|\mathbf{S}^{(K)}_{i}-\mathbf{c}_{\hat{\sigma}(i)}\right\|_{2}^{2}.$ (49)
-Recall the definition of $\bar{\sigma}(\cdot)$ in (1) we have $\bar{\sigma}(\pi(i))=\hat{\sigma}(i)$ for some permutation $\pi(\cdot)$.
+Recall the definition of $\bar{\sigma}(\cdot)$ in (1) we have $\pi(\bar{\sigma}(i))=\hat{\sigma}(i)$ for some permutation $\pi(\cdot)$.
 Proof.